### v1

# Glivenko–Cantelli for $f$-divergence

Abstract.
We extend the celebrated Glivenko–Cantelli theorem, sometimes called the fundamental theorem of statistics, from its standard setting of total variation distance to all $f$-divergences. A key obstacle in this endeavor is to define $f$-divergence on a subcollection of a $\sigma$-algebra that forms a $\pi$-system but not a $\sigma$-subalgebra. This is a side contribution of our work. We will show that this notion of $f$-divergence on the $\pi$-system of rays preserves nearly all known properties of standard $f$-divergence, yields a novel integral representation of the Kolmogorov–Smirnov distance, and has a Glivenko–Cantelli theorem.
Key words and phrases:
probability measures, Kolmogorov–Smirnov distance, total variation distance, $f$-divergence, Glivenko–Cantelli theorem
1991 Mathematics Subject Classification:
28A33, 28A50, 46E27, 60E05, 60E15, 60F17, 94A17
1. Introduction
The Glivenko–Cantelli theorem [22, 8] is a cornerstone of empirical process theory. It is likely the best known statement regarding the asymptotic behavior of stochastic processes formed by empirical measures [16]. It is also crucial in nonparametric statistics and forms the basis for statistical consistency in many estimation problems [16]. In the Kolmogorov–Smirnov test, the theorem ensures that the test statistic has desirable asymptotic properties [46]. In statistical resampling methods like the bootstrap, it guarantees that the empirical distribution derived from resampled data will approximate the true distribution as the number of samples grows [18]. Because of its many central roles, the Glivenko–Cantelli theorem is often called the fundamental theorem of statistics [49, 27, 14, 44].
In machine learning, the Glivenko–Cantelli theorem can be extended to the Vapnik–Chervonenkis theorem [55], also known as the fundamental theorem of learning theory [45]. It is used to show the consistency of the principle of empirical risk minimization (ERM) [54] and to derive bounds on generalization error by ensuring the difference between empirical and true distributions diminishes with larger samples, aiding in over-fitting control [6].
Given its pivotal role in both statistics and machine learning, it is surprising that the Glivenko–Cantelli theorem is somewhat limited in scope—it only works with the total variation norm restricted to a Glivenko–Cantelli class. Modern AI applications, on the other hand, have ushered in a plethora of alternative distances between two probability distributions $\mu$ and $\nu$, most of them easier to compute than the total variation norm. The most prominent of these are the $f$-divergences
(1) $\operatorname{D}_{f}(\mu\|\nu)=\int_{\Omega}f\Bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu=\int_{\Omega}f\bigl{(}g(x)\bigr{)}\,\mathrm{d}\nu(x),$
where $g$ is the Radon–Nikodym derivative with $\mathrm{d}\mu(x)=g(x)\,\mathrm{d}\nu(x)$. The total variation norm is itself an $f$-divergence with $f(t)=\lvert t-1\rvert/2$ but many other $f$-divergence with yet other $f$’s have played a prominent role in important AI applications recently. In [31], the Kullback–Leibler (KL) divergence is used to regularize the posterior distribution in the Variational Autoencoder (VAE), aligning it with a chosen prior. In [23], Generative Adversarial Networks (GANs) are trained by minimizing the Jensen–Shannon (JS) divergence between the model and real data distributions. In [33], Rényi $\alpha$-divergences provide a smooth interpolation from the evidence lower bound to the log marginal likelihood that is controlled by the parameter $\alpha$. The goal of our article is to generalize the Glivenko–Cantelli theorem to any $f$-divergence, restricted to a Glivenko–Cantelli class defined with respect to this $f$-divergence.
At this point, one may wonder about Wasserstein metrics, another class of distances between probability distributions that rivals $f$-divergences in importance and popularity. But as was shown in [5, 35], there is no equivalent of Glivenko–Cantelli theorem for $p$-Wasserstein metric. Such a result will always be trivial as the “Glivenko–Cantelli class” with respect to any $p$-Wasserstein metric is just the entire Borel $\sigma$-algebra: see [5] for $p\in[1,\infty)$ and [35] for $p=\infty$. On the other hand, the main contribution of our article shows that there is a natural, nontrivial generalization of the Glivenko–Cantelli theorem to $f$-divergences of which the special case $f(t)=\lvert t-1\rvert/2$ gives us the classical Glivenko–Cantelli theorem.
We remind the reader of this classical version, taking the opportunity to set notations, define basic notions, and establish some facts for later use. Let $(\Omega,\Sigma,P)$ be a probability space and $X:\Omega\to\mathbb{R}$ be a random variable. Let $\nu$ be the distribution of $X$, i.e., the pushforward measure of $P$ by $X$. Thus $\nu$ is a probability measure on the real Borel space $(\mathbb{R},\mathcal{B})$. For $X_{1},\dots,X_{n}\sim\nu$ independent and identically distributed random variables on the probability space $(\mathbb{R},\mathcal{B},\nu)$, the empirical distribution is the random measure
$\nu_{n}(A,\omega)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}_{\{X_{i}(\omega)\in A\}},$
defined for any $A\in\mathcal{B}$ and $\omega\in\Omega$. For a fixed $\omega\in\Omega$, $\nu_{n}(\cdot,\omega)$ is a probability measure on $(\mathbb{R},\mathcal{B})$; and for a fixed $A\in\mathcal{B}$, $\nu_{n}(A)\coloneqq\nu_{n}(A,\cdot)$ is a random variable on $\Omega$.
As is well known, for any fixed $A\in\mathcal{B}$, the strong law of large number states that $\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}\to 0$ almost surely as $n\to\infty$, i.e.,
SLLN: for each fixed $\text{for each fixed }A,\;\nu_{n}(A)-\nu(A)\to 0\text{ with probability one.}$.
The Glivenko–Cantelli theorem may be viewed as an attempt to replace the qualifier “for each fixed $A$” by “for all $A\in\mathcal{C}$” where $\mathcal{C}$ is an appropriately chosen class of subsets of $\Omega$ called a Glivenko–Cantelli class [46], while maintaining the property that as $n\to\infty$,
GC: $\sup_{A\in\mathcal{C}}\,\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}\to 0\text{ with probability one.}$.
The following example shows that one cannot simply take $\mathcal{C}=\mathcal{B}$, the Borel $\sigma$-algebra.
Example 1.1.
Let $\nu$ be the standard Gaussian distribution. Let $X_{1},\dots,X_{n}\sim\nu$ be independent and identically distributed random variables with realizations $x_{1},\dots,x_{n}$. If $A=\{x_{1},\dots,x_{n}\}\in\mathcal{B}$, then $\nu_{n}(A)=1$ but $\nu(A)=0$. Thus $\sup_{A\in\mathcal{B}}\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}=1$ for all $n$ and $\mathcal{B}$ is not a Glivenko–Cantelli class.
One might think that $\mathcal{B}$ is too big but even if we choose a much smaller $\mathcal{C}$, it could still fail to be a Glivenko–Cantelli class, as the following example from [47] shows.
Example 1.2 (Shortt).
Let $\nu$ be the uniform measure on $[0,1]$ and $\mathcal{C}$ be the collection of all finite subset of $[0,1]$. Let $X_{1},\dots,X_{n}\sim\nu$ be independent and identically distributed random variables with realizations $x_{1},\dots,x_{n}$. If $A=\{x_{1},\dots,x_{n}\}\in\mathcal{C}$, then $\nu_{n}(C)=1$ but $\nu(C)=0$. Thus $\sup_{A\in\mathcal{C}}\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}=1$ for all $n$ and $\mathcal{C}$ is not a Glivenko–Cantelli class.
So the existence of (nontrivial) Glivenko–Cantelli classes is not obvious. The Glivenko–Cantelli theorem shows that they indeed exists, with the class of rays the original [22, 8] and best known example.
Definition 1.3 (Rays).
The class of open rays and the class of closed rays are
$\mathcal{R}_{\mathrm{o}}=\{(-\infty,a):a\in\mathbb{R}\}\cup\{\varnothing,\mathbb{R}\},\quad\mathcal{R}_{\mathrm{c}}=\{(-\infty,a]:a\in\mathbb{R}\}\cup\{\varnothing,\mathbb{R}\}$
respectively. Their union is the class of rays, denoted
$\mathcal{R}=\mathcal{R}_{\mathrm{c}}\cup\mathcal{R}_{\mathrm{o}}.$
These are all $\pi$-systems [17, p. 202] but not $\sigma$-algebras. In this article, we restrict ourselves to “left” rays, i.e., where the left-end point of the interval is always $\infty$- ∞. In topology, $\mathcal{R}_{\mathrm{o}}$ is called the left-ray topology or left-order topology.
Glivenko [22] showed that $\mathcal{R}_{\mathrm{c}}$ is a Glivenko–Cantelli class whereas Cantelli [8] showed that $\mathcal{R}_{\mathrm{o}}$ is a Glivenko–Cantelli class but we will soon see that it makes no difference whether we use $\mathcal{R}_{\mathrm{c}}$, $\mathcal{R}_{\mathrm{o}}$, or $\mathcal{R}$; all three statements are equivalent. Obviously, the name “Glivenko–Cantelli class” came about much later [46], earlier statements of the result were more of the following form:
Theorem 1.4 (Glivenko–Cantelli).
Let $\nu$ be a Borel probability measure and $\nu_{n}$ be the corresponding empirical measure, $n\in\mathbb{N}$. Then, almost surely,
(2) $\sup_{A\in\mathcal{R}}\,\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}\to 0.$
There is no substantive difference if we instead use $\mathcal{R}_{\mathrm{o}}$ or $\mathcal{R}_{\mathrm{c}}$ in place of $\mathcal{R}$ in Theorem 1.4 as
(3) $\sup_{A\in\mathcal{R}_{\mathrm{c}}}\bigl{(}\mu(A)-\nu(A)\bigr{)}=\sup_{A\in\mathcal{R}_{\mathrm{o}}}\bigl{(}\mu(A)-\nu(A)\bigr{)}=\sup_{A\in\mathcal{R}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}$
for any probability measures $\mu$ and $\nu$ on $(\mathbb{R},\mathcal{B})$, a fact that follows easily from the continuity of probability measures. The common value in (3) defines a distance between probability measures called the Kolmogorov–Smirnov distance, although it is usually defined in terms of $\mathcal{R}_{\mathrm{c}}$ [30].
The total variation distance is defined by
(4) $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)=\sup_{A\in\mathcal{B}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}.$
So the Kolmogorov–Smirnov distance is just the total variation distance “restricted” to $\mathcal{R}\subseteq\mathcal{B}$, i.e., a partial variation over a smaller class of sets than the full Borel $\sigma$-algebra. As we have alluded to earlier, the total variation is an example of $f$-divergence, an extensive class of distance between probability measures. Note that having absolute values in (4) makes no difference because $A\in\mathcal{B}$ iff $A^{\mathsf{c}}\in\mathcal{B}$. Throughout this article, an integral sign $\int$ unadorned with upper and lower limits is taken to mean $\int_{\mathbb{R}}$.
Definition 1.5 ($f$-divergence).
Let $f:\mathbb{R}\to\mathbb{R}$ be a continuous convex function with $f(1)=0$. Let $\mu$ and $\nu$ be probability measures on $(\mathbb{R},\mathcal{B})$ be such that $\mu\ll\nu$, i.e., $\mu$ is absolutely continuous with respect to $\nu$, so that we may speak of Radon–Nikodym derivative $\mathrm{d}\mu/\mathrm{d}\nu$. Then the $f$-divergence [41, 13, 37, 1] between $\mu$ and $\nu$ is given by
(5) $\operatorname{D}_{f}(\mu\|\nu)=\int f\Bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu.$
Different choices of $f$ yield various well-known divergences, including Kullback–Liebler [3], Le Cam [32], Jensen–Shannon [34], Jeffreys [29], Chernoff [10], Pearson $\chi^{2}$ [39], Hellinger squared [26], exponential [36], and alpha–beta [19] divergences, and so on. For $f(t)=\lvert t-1\rvert/2$, we get [52]:
$\int\frac{1}{2}\Bigl{\lvert}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{\rvert}\,\mathrm{d}\nu=\sup_{A\in\mathcal{B}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)},$
the total variation distance. To obtain a Glivenko–Cantelli theorem for an arbitrary $f$-divergence, we first need an analog of Kolmogorov–Smirnov distance corresponding to an $f$-divergence.
Question 1.
Is there a notion of $f$-divergence over $\mathcal{R}$ for general $f$ that (a) reduces to the Kolmogorov–Smirnov distance in (3) when $f(t)=\lvert t-1\rvert/2$; and (b) reduces to the standard $f$-divergence in (5) when $\mathcal{R}$ is replaced by $\mathcal{B}$?
The answer to Question 1 is yes, provided by what we will call an $f$-divergence over $\mathcal{R}$ and denoted $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)$. We will establish the existence of $f$-divergence over $\mathcal{R}$ in Section 3 and develop some of its properties in Section 4, showing in particular that this new notion preserves some of the best-known relations between different $f$-divergences. In Section 5, we will prove the result claimed in the title of our article, namely, $\operatorname{D}_{f}^{\mathcal{R}}(\nu_{n}\|\nu)$ converges to zero almost surely. Note that when $f(t)=\lvert t-1\rvert/2$, this reduces to the Glivenko–Cantelli theorem in Theorem 1.4.
We emphasize that the groundbreaking work of Vapnik and Chervonenkis [55] sheds no light on whether the Glivenko–Cantelli theorem remains true when total variation distance is replaced by an $f$-divergence. They showed that if the VC dimension of a class $\mathcal{C}\subseteq\Sigma$, i.e., the maximum number of points that $\mathcal{C}$ can shatter, is finite, then $\mathcal{C}$ is a Glivenko–Cantelli class, i.e., $\sup_{A\in\mathcal{C}}\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}\to 0$ almost surely. Evidently, the Vapnik–Chervonenkis theorem is still entirely about the total variation distance.
As we will see, it takes some effort to establish the notion of $f$-divergence over $\mathcal{R}$. For the total variation distance, getting from (4) to (3) is a matter of simply replacing $\mathcal{B}$ by $\mathcal{R}$. But for general $f$-divergence, $\mathcal{B}$ does not even appear directly in (5) and it is no longer a matter of simply replacing $\mathcal{B}$ by $\mathcal{R}$. Indeed we know of no straightforward answer to Question 1. Readers familiar with Choquet integrals [11, 15] may think that it provides a plausible answer, as we did initially; we will explain why it does not in Section 6.1. We will discuss the prospects of a “Vapnik–Chervonenkis theorem for $f$-divergence” in Section 6.2.
2. Notations and terminologies
2.1. Functions
We reserve the letter $f$ for $f$-divergence in this article. Other functions will be denoted $g$, $h$, $\varphi$, $\psi$. Depending on context, the notation $X_{k}\uparrow X$ could mean an increasing sequence $X_{1}\leq X_{2}\leq\dots\leq X_{k}\in\mathbb{R}$ converging to a real value $X\in\mathbb{R}$ or a nested sequence of sets $X_{1}\subseteq X_{2}\subseteq\dots\subseteq X_{k}\subseteq\mathbb{R}$ converging to a set $X\subseteq\mathbb{R}$. Similarly for $X_{k}\downarrow X$.
[…]
$L^{2}_{+}(\nu)\coloneqq\{g\in L^{2}(\nu):g\geq 0\}$
for the cone of functions nonnegative almost everywhere. As usual, by “function” we mean an equivalence class of functions that differ at most on a measure-zero set and conditions like “$h\geq g$” will always be in the almost everywhere sense without specification. On the other hand, we will specify the measure every time since we often have to deal with multiple measures.
[…]
Let $\varnothing\neq M\subseteq L^{2}(\nu)$ be a closed convex subset. The metric projection onto $M$ is the operator $\operatorname{proj}_{M}:L^{2}(\nu)\to L^{2}(\nu)$ that takes $g\in L^{2}(\nu)$ to the closest point $g_{*}\in M$, i.e., $\lVert g-g_{*}\rVert=\min_{h\in M}\lVert h-g\rVert$. The existence and uniqueness of $g_{*}$ is guaranteed by the conditions on $M$ and so $\operatorname{proj}_{M}$ is well-defined. It is also well-known that $\operatorname{proj}_{M}$ is continuous [2, p. 52].
2.2. Sequences
We will remind the readers of the three notions of convergence used in our article: For a sequence of measurable functions $g_{n}$ converging to $g$ in $L^{2}(\nu)$. The convergence is said to be (i) $\nu$-almost surely if $g_{n}$ converges to $g$ pointwise except on a $\nu$-null set; (ii) in $\nu$-measure if $\lim_{n\to\infty}\nu(\lvert g_{n}-g\rvert>\varepsilon)=0$ for all $\varepsilon>0$; (iii) in $L^{2}$-norm if $\lVert g_{n}-g\rVert\to 0$. The following lemma, adapted from [12, pp. 80–83], summarizes the relations between them.
Lemma 2.1.
Let $\nu$ be a Borel probability measure.
[…]
    If $g_{n}\to g$ in $L^{2}$-norm, then $g_{n}\to g$ in $\nu$-measure.
[…]
    If $g_{n}\to g$ in $\nu$-measure, then there is a subsequence $(g_{n_{k}})$ such that $g_{n_{k}}\to g$ almost surely.
[…]
    If $g_{n}\to g$ almost surely and there exists $g\in L^{2}(\nu)$ that dominates $g_{n}$, then $g_{n}\to g$ in $L^{2}$-norm.
We will also need an observation that follows from $\lvert\int g_{n}-g\,\mathrm{d}\nu\rvert\leq\int\lvert g_{n}-g\rvert\,\mathrm{d}\nu\leq\bigl{(}\int(g_{n}-g)^{p}\,\mathrm{d}\nu\bigr{)}^{1/p}$.
Lemma 2.2.
Let $g_{n}\in L^{p}(\nu)$. If $g_{n}\to g$ in $L^{p}$-norm for some $1\leq p<\infty$, then
$\int g\,\mathrm{d}\nu=\lim_{n\to\infty}\int g_{n}\,\mathrm{d}\nu.$
2.3. Measures
We write $2^{\Omega}$ for the power set of $\Omega$. A collection of subsets $\mathcal{E}\subseteq 2^{\Omega}$ is called a $\pi$-system if it is closed under finite intersections [17, p. 202]. The following results from [12, p. 38] and [43, p. 58] are reproduced here for easy reference.
Lemma 2.3.
Let $\mu,\nu$ be finite measures on $(\Omega,\Sigma)$ such that $\mu(\Omega)=\nu(\Omega)$. If $\mathcal{A}$ is a $\pi$-system that generates $\Sigma$ and $\mu(A)=\nu(A)$ for any $A\in\mathcal{A}$, then $\mu=\nu$.
Lemma 2.4.
Let $g:X\to Y$ be any function and $\mathcal{E}\subseteq 2^{Y}$. Then $\sigma(g^{-1}(\mathcal{E}))=g^{-1}(\sigma(\mathcal{E}))$.
Whenever we write $\mathrm{d}\mu/\mathrm{d}\nu$, we implicitly assume that $\mu\ll\nu$.
3. $f$-divergence over rays
The goal of this section is to resolve Question 1 in the affirmative. The answer is an analogue of $f$-divergence with respect to the class of rays $\mathcal{R}$ that we will call the $f$-divergence over $\mathcal{R}$. This appears only at the end of this section in Definition 3.12. It will take the groundwork developed over the course of the whole section before we can show that the notion is indeed well-defined. There are two milestones to watch for: We will see in Theorem 3.15 that the construction of such a divergence would work for any class $\mathcal{C}\subseteq\mathcal{B}$ that satisfies a certain “Radon–Nikodym property.” We will see in Corollary 3.11, which follows from Theorem 3.9, that $\mathcal{C}=\mathcal{R}$ has this property.
We begin by establishing some basic properties related to $\mathcal{R}$. Throughout this section, $\nu$ denotes a Borel probability measure on $(\mathbb{R},\mathcal{B})$.
Proposition 3.1.
The set
$G(\mathcal{R})\coloneqq\bigl{\{}g\in L^{2}(\nu):\{g>r\}\in\mathcal{R}\text{ for all }r\in\mathbb{R}\bigr{\}}$
is a closed convex cone in $L^{2}(\nu)$ comprising all nonincreasing functions in $L^{2}(\nu)$.
Proof.
We first show that $G(\mathcal{R})$ is exactly the set of nonincreasing $L^{2}$-functions. Let $g$ be such a function and let $r\in\mathbb{R}$. Set $a=\sup\{x:g(x)>r\}$. Any $x\in\{g>r\}$ has $x\leq a$ so $\{g>r\}\subseteq(-\infty,a]$. If there is an $x<a$ with $g(x)\leq r$, then there must be some $y>x$ with $g(y)>r\geq g(x)$. This contradicts the assumption that $g$ is nonincreasing. Hence $(\infty,a)\subseteq\{g>r\}\subseteq(-\infty,a]$. It follows that $\{g>r\}\in\mathcal{R}$ and so $g\in G(\mathcal{R})$. Conversely, let $g\in G(\mathcal{R})$. Suppose we have $x<y$ with $g(x)<g(y)$. Then there is some $a$ with $(-\infty,a)\subseteq\{g>g(x)\}\subseteq(-\infty,a]$. Since $g(y)>g(x)$, it follows that $y\in\{g>g(x)\}$, and so $x<y<a$. Consequently, $x\in\{g>g(x)\}$, i.e., $g(x)>g(x)$, a contradiction. Hence $g$ must be nonincreasing.
$G(\mathcal{R})$ is a closed convex cone because nonnegative linear combinations or $L^{2}$-limits of nonincreasing functions remain nonincreasing. The former is routine. The latter follows from Lemma 2.1: A sequence $(g_{n})$ of nonincreasing functions with $\lVert g_{n}-g\rVert\to 0$ must also converge in $\nu$-measure. So there is a subsequence $(g_{n_{k}})$ that converges to $g$ almost surely. So there is a null set $A$ where $\lim_{k\to\infty}g_{n_{k}}(x)=g(x)$ for all $x\notin A$. Since each $g_{n_{k}}$ is nonincreasing, $g$ is almost surely nonincreasing. ∎
An immediate consequence of Proposition 3.1 is that $G(\mathcal{R})$ has a well-defined metric projection. For a given $g\in L^{2}(\nu)$, finding $\operatorname{proj}_{G(\mathcal{R})}g$ is equivalent to finding the nearest nonincreasing function to $g$ in $L^{2}$-norm.
Definition 3.2.
A collection $\{A_{1},\dots,A_{n}\}$ is called an ordered partition of $\mathbb{R}$ if it satisfies the following conditions:
[…]
    disjoint: $A_{i}\cap A_{j}=\varnothing$ for any $i\neq j$;
[…]
    cover: $A_{1}\cup\dots\cup A_{n}=\mathbb{R}$;
[…]
    ordered: $\sup A_{i}\leq\inf A_{j}$ for each $i\leq j$.
An ordered simple function $h:\mathbb{R}\to\mathbb{R}$ is a simple function defined on an ordered partition $\{A_{1},\dots,A_{n}\}$.
It is well-known that any $g\in L^{2}_{+}(\nu)$ can be approximated to arbitrary accuracy by an increasing sequence of nonnegative simple functions $(h_{n})$. We will show in Corollary 3.4 that $h_{n}$ can be chosen to be ordered simple functions.
Proposition 3.3.
If $g\in L^{2}_{+}(\nu)$ is almost surely continuous, then there exists a monotone increasing sequence of nonnegative ordered simple functions $(h_{n})$ that converges pointwise to $g$ almost surely.
Proof.
Define the ordered partition
$\mathcal{P}_{n}\coloneqq\Bigl{\{}\Bigl{[}\frac{nk}{2^{n}},\frac{n(k+1)}{2^{n}}\Bigr{)}:k=-2^{n},-2^{n}+1,\dots,2^{n}-1\Bigr{\}}\cup\bigl{\{}(-\infty,-n),[n,\infty)\bigr{\}}$
and the ordered simple function $h_{n}=\sum_{A\in\mathcal{P}_{n}}\inf_{x\in A}g(x)\mathbbm{1}_{A}$. Since $g$ is almost surely continuous, we can find a null set $S$ such that for any $\varepsilon>0$ and $\𝑆x\in\mathbb{R}\backslash S$, there exist $n\in\mathbb{N}$ and $A\in\mathcal{P}_{n}$ with $x\in A\subseteq B_{\varepsilon}(x)$. Consequently,
$\inf_{y\in B_{\varepsilon}(x)}g(y)\leq\inf_{y\in A}g(y)=h_{n}(x)\leq g(x),$
[…]
In Proposition 3.3, $h_{n}$ is dominated by $g$; so we also have $h_{n}\to g$ in $L^{2}$-norm by Lemma 2.1. The almost sure continuity in Proposition 3.3 may sometimes be overly restrictive. We derive an alternative version without this requirement.
Corollary 3.4.
Let $g\in L^{2}_{+}(\nu)$. Then there exists a sequence of bounded ordered simple functions $(h_{n})$ that converges to $g$ in $L^{2}$-norm and almost surely.
Proof.
Given that the space of continuous functions with compact support $C_{\mathrm{c}}(\mathbb{R})$ is dense in $L^{2}(\nu)$ [28, p. 354], there is a continuous function $\varphi$ with compact support that approximates $g$ to accuracy $\lVert g-\varphi\rVert<\varepsilon/2$ for any $\varepsilon>0$. On the other hand, the continuous function $\varphi$ is in $L^{2}(\nu)$; so by Proposition 3.3 there is an ordered simple function $h$ that approximates it to accuracy $\lVert\varphi-h\rVert<\varepsilon/2$. Hence $\lVert h-g\rVert<\varepsilon$. This shows that for any $g\in L^{2}_{+}(\nu)$, there exists a sequence of ordered simple function $h_{n}$ such that $h_{n}\to g$ in $L^{2}$-norm. By passing through a subsequence if necessary, Lemma 2.1 ensures $h_{n}\to g$ almost surely as well. ∎
We next see that the metric projection of an ordered simple function onto $G(\mathcal{R})$ remains ordered simple.
Proposition 3.5 (Projection).
Let $h=\sum_{i=1}^{n}\alpha_{i}\mathbbm{1}_{A_{i}}\in L^{2}_{+}(\nu)$ be an ordered simple function with $\{A_{1},\dots,A_{n}\}$ an ordered partition of $\mathbb{R}$. Then $\operatorname{proj}_{G(\mathcal{R})}h$ is also an ordered simple function on $\{A_{1},\dots,A_{n}\}$.
Proof.
Suppose that $\operatorname{proj}_{G(\mathcal{R})}h$ is not constant on $A_{j}$ for some $j=1,\dots,n$. Define
$\varphi=\begin{cases}\operatorname{proj}_{G(\mathcal{R})}h&\text{on }A_{j}^{\mathsf{c}},\\ c&\text{on }A_{j},\end{cases}$
where $c=\operatorname{proj}_{G(\mathcal{R})}h(\omega_{*})$ and $\omega_{*}=\operatorname*{argmin}_{\omega\in A_{j}}\lvert h(\omega)-\operatorname{proj}_{G(\mathcal{R})}h(\omega)\rvert$. Then $\varphi\in G(\mathcal{R})$ and
$\int(h-\operatorname{proj}_{G(\mathcal{R})}h)^{2}\,\mathrm{d}\nu\geq\int(h-\varphi)^{2}\,\mathrm{d}\nu,$
a contradiction. ∎
A consequence of Proposition 3.5 is that metric projection onto $G(\mathcal{R})$ preserves ordering.
Proposition 3.6 (Monotonicity).
Let $g_{1},g_{2}\in L^{2}_{+}(\nu)$. If $g_{1}\leq g_{2}$, then
$\operatorname{proj}_{G(\mathcal{R})}g_{1}\leq\operatorname{proj}_{G(\mathcal{R})}g_{2}.$
Proof.
We first show that the statement holds for ordered simple functions. Let $h_{1}=\sum_{i=1}^{n}a_{i}\mathbbm{1}_{A_{i}}$ and $h_{2}=\sum_{j=1}^{m}b_{j}\mathbbm{1}_{B_{j}}$ where $(A_{1},\dots,A_{n})$ and $(B_{1},\dots,B_{m})$ are ordered partitions. Suppose $h_{1}\leq h_{2}$ but that $\operatorname{proj}_{G(\mathcal{R})}h_{1}>\operatorname{proj}_{G(\mathcal{R})}h_{2}$ on $E=A_{i}\cap B_{j}$ for some $i$ and $j$. Define
$\varphi$ $=(\operatorname{proj}_{G(\mathcal{R})}h_{1})\cdot\mathbbm{1}_{E^{\mathsf{c}}}+(\operatorname{proj}_{G(\mathcal{R})}h_{2})\cdot\mathbbm{1}_{E},$
$\psi$ $=(\operatorname{proj}_{G(\mathcal{R})}h_{2})\cdot\mathbbm{1}_{E^{\mathsf{c}}}+(\operatorname{proj}_{G(\mathcal{R})}h_{1})\cdot\mathbbm{1}_{E}.$
Both $\varphi$ and $\psi$ are clearly nonincreasing and thus $\mathcal{R}$-measurable. There are six possibilities for the values of these functions on $E$:
[…]
    $h_{2}\geq h_{1}\geq\operatorname{proj}_{G(\mathcal{R})}h_{1}\geq\operatorname{proj}_{G(\mathcal{R})}h_{2}$,
[…]
    $h_{2}\geq\operatorname{proj}_{G(\mathcal{R})}h_{1}\geq h_{1}\geq\operatorname{proj}_{G(\mathcal{R})}h_{2}$,
[…]
    $h_{2}\geq\operatorname{proj}_{G(\mathcal{R})}h_{1}\geq\operatorname{proj}_{G(\mathcal{R})}h_{2}\geq h_{1}$,
[…]
    $\operatorname{proj}_{G(\mathcal{R})}h_{1}\geq h_{2}\geq h_{1}\geq\operatorname{proj}_{G(\mathcal{R})}h_{2}$,
[…]
    $\operatorname{proj}_{G(\mathcal{R})}h_{1}\geq h_{2}\geq\operatorname{proj}_{G(\mathcal{R})}h_{2}\geq h_{1}$,
[…]
    $\operatorname{proj}_{G(\mathcal{R})}h_{1}\geq\operatorname{proj}_{G(\mathcal{R})}h_{2}\geq h_{2}\geq h_{1}$.
For Cases i, ii, and iii,
$\lVert\psi-h_{2}\rVert\leq\lVert\operatorname{proj}_{G(\mathcal{R})}h_{2}-h_{2}\rVert;$
for Cases v and vi,
$\lVert\varphi-h_{1}\rVert\leq\lVert\operatorname{proj}_{G(\mathcal{R})}h_{1}-h_{1}\rVert;$
and for Case iv, either of the following
$\lVert\psi-h_{2}\rVert\leq\lVert\operatorname{proj}_{G(\mathcal{R})}h_{2}-h_{2}\rVert,\quad\lVert\varphi-h_{1}\rVert\leq\lVert\operatorname{proj}_{G(\mathcal{R})}h_{1}-h_{1}\rVert$
must hold; all leading to contradictions.
For the general case with $g_{1},g_{2}\in L^{2}_{+}(\nu)$, we set $g_{0}\coloneqq g_{2}-g_{1}\in L^{2}_{+}(\nu)$. By Corollary 3.4, we have two sequences of nonnegative ordered simple functions $(h_{0,n})$ and $(h_{1,n})$ such that $h_{0,n}\to g_{0}$ and $h_{1,n}\to g_{1}$ in $L^{2}$-norm. Set $h_{2,n}\coloneqq h_{0,n}+h_{1,n}$, which is nonnegative, ordered simple, and satisfies $h_{2,n}\to g_{2}$ in $L^{2}$-norm and $h_{2,n}\geq h_{1,n}$ for each $n$. By the continuity of the metric projection operator $\operatorname{proj}_{G(\mathcal{R})}$,
$\operatorname{proj}_{G(\mathcal{R})}g_{1}=\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}h_{1,n}\leq\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}h_{2,n}=\operatorname{proj}_{G(\mathcal{R})}g_{2},$
as required. ∎
A slight refinement of the last step of the proof above yields the following.
Corollary 3.7 (Sequential monotonicity).
Let $g\in L^{2}_{+}(\nu)$. If $(h_{n})$ is a monotone increasing sequence in $L^{2}_{+}(\nu)$ converging to $g$, then $(\operatorname{proj}_{G(\mathcal{R})}h_{n})$ is a monotone increasing sequence in $L^{2}_{+}(\nu)$ converging to $\operatorname{proj}_{G(\mathcal{R})}g$. Convergence is both in the sense of $L^{2}$-norm and almost surely.
Proof.
Let $h_{n}$ be a sequence of ordered simple function such that $h_{n}\to g$ almost surely and in $L^{2}$-norm. Continuity of $\operatorname{proj}_{G(\mathcal{R})}$ yields
$\operatorname{proj}_{G(\mathcal{R})}h_{n}\to\operatorname{proj}_{G(\mathcal{R})}g$
in $L^{2}$-norm. By passing to a subsequence if necessary, we may assume that this convergence also holds in the almost sure sense. ∎
Proposition 3.8.
Let $A\in\mathcal{R}$. Then
$\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})=\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})\cdot\mathbbm{1}_{A}\leq(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{A}.$
Proof.
We have
$\int(g\mathbbm{1}_{A}-$ $\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A}))^{2}\,\mathrm{d}\nu$
$=\int_{A}(g\mathbbm{1}_{A}-\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A}))^{2}\,\mathrm{d}\nu+\int_{A^{\mathsf{c}}}(g\mathbbm{1}_{A}-\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A}))^{2}\,\mathrm{d}\nu$
$=\int_{A}(g\mathbbm{1}_{A}-\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A}))^{2}\,\mathrm{d}\nu+\int_{A^{\mathsf{c}}}(\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A}))^{2}\,\mathrm{d}\nu$
$\geq\int_{A}(g\mathbbm{1}_{A}-\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A}))^{2}\,\mathrm{d}\nu$
$=\int(g\mathbbm{1}_{A}-\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})\cdot\mathbbm{1}_{A})^{2}\,\mathrm{d}\nu.$
Since the projection is unique, we have $\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})=\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})\cdot\mathbbm{1}_{A}$. As $g\mathbbm{1}_{A}\leq g$, the monotonicity in Proposition 3.6 yields
$\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})=\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})\cdot\mathbbm{1}_{A}\leq(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{A}.\qed$
Let $\mathcal{C}\subseteq\mathcal{B}$ be a $\sigma$-subalgebra. The Radon–Nikodym theorem states that for $\sigma$-finite measures $\mu$ and $\nu$ on the measurable space $(\mathbb{R},\mathcal{B})$ with $\mu\ll\nu$, there exists a $\mathcal{C}$-measurable function $\rho_{\mathcal{C}}$ such that $\mu(A)=\int_{A}\rho_{\mathcal{C}}\,\mathrm{d}\nu$ for any $A\in\mathcal{C}$.
What happens if $\mathcal{C}$ is not a $\sigma$-subalgebra? For us, the most important case to consider is when $\mathcal{C}=\mathcal{R}$, which is not a $\sigma$-subalgebra. We will show in Corollary 3.11 that $\mathcal{R}$ satisfies a kind of Radon–Nikodym property: There is a $\mathcal{R}$-measurable function $\rho_{\mathcal{R}}$ and a subcollection $\mathcal{E}\subseteq\mathcal{R}$ such that $\mu(A)=\int_{A}\rho_{\mathcal{R}}\,\mathrm{d}\nu$ for any $A\in\mathcal{E}$. The bulk of the work is in establishing the following technical result.
Theorem 3.9.
Let $g\in L^{2}_{+}(\nu)$. Define the $\pi$-system $\mathcal{E}(g)\subseteq\mathcal{R}$ by
(6) $\mathcal{E}(g)\coloneqq\bigl{\{}\{\operatorname{proj}_{G(\mathcal{R})}g>r\}:r\geq 0\bigr{\}}\cup\{\mathbb{R}\}.$
Then for any $E\in\mathcal{E}(g)$,
(7) $\int_{E}\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu=\int_{E}g\,\mathrm{d}\nu.$
Proof.
We establish the result in four steps, showing that
[…]
    it holds when $g$ is an ordered simple function $h$;
[…]
    it holds for $E=\{\operatorname{proj}_{G(\mathcal{R})}g>r\}$ under the assumption that $\nu(\{\operatorname{proj}_{G(\mathcal{R})}g=r\})=0$;
[…]
    it extends to all $E=\{\operatorname{proj}_{G(\mathcal{R})}g>r\}$;
[…]
    it holds for $E=\mathbb{R}$.
Step i: First we assume that we have an ordered simple function
$h=\sum_{i=1}^{n}\alpha_{i}\mathbbm{1}_{A_{i}},\quad\operatorname{proj}_{G(\mathcal{R})}h=\sum_{i=1}^{n}\beta_{i}\mathbbm{1}_{A_{i}}.$
So $E=\{\operatorname{proj}_{G(\mathcal{R})}h>r\}$ can be written as $E=A_{1}\cup\dots\cup A_{k}\in\mathcal{E}(h)$ where $\beta_{k}>r$ and $\beta_{k+1}\leq r$. Write $\alpha=(\alpha_{1},\dots,\alpha_{n})$, $\beta=(\beta_{1},\dots,\beta_{n})$, $\omega=(\omega_{1},\dots,\omega_{n})$ where $\omega_{i}=\nu(A_{i})$, and let $W=\mathrm{diag}(\omega)$. Then $\beta$ is the solution to the following quadratic programming problem:
$\operatorname{minimize}$ $\quad(\alpha-\beta)^{\scriptscriptstyle\mathsf{T}}W(\alpha-\beta)$
$\operatorname{subject~{}to}$ $\quad\beta_{1}\geq\dots\geq\beta_{n}\geq 0.$
Observe that for
$\Pi=\begin{bmatrix}1&-1&&\\ &1&-1&\\ &&\ddots&-1\\ &&&1\end{bmatrix},\qquad\Pi^{-1}=\begin{bmatrix}1&1&\dots&1\\ &1&\dots&1\\ &&\ddots&\vdots\\ &&&1\end{bmatrix}.$
Let $\gamma=\Pi\beta$. Then the quadratic programming problem above is equivalent to
$\operatorname{minimize}$ $\quad\gamma^{\scriptscriptstyle\mathsf{T}}\Pi^{-{\scriptscriptstyle\mathsf{T}}}W\Pi^{-1}\gamma-2\alpha^{\scriptscriptstyle\mathsf{T}}\Pi^{-1}W\gamma$
$\operatorname{subject~{}to}$ $\quad-\gamma\preceq 0.$
Standard KKT conditions yield
$\lambda=2\Pi^{-{\scriptscriptstyle\mathsf{T}}}W(\Pi^{-1}\gamma-\alpha)=2\Pi^{-{\scriptscriptstyle\mathsf{T}}}W(\beta-\alpha)$
or
(8) $\begin{bmatrix}\lambda_{1}\\ \lambda_{2}\\ \vdots\\ \lambda_{n}\end{bmatrix}=2\begin{bmatrix}1&&&\\ 1&1&&\\ \vdots&\vdots&\ddots&\\ 1&1&\dots&1\end{bmatrix}\begin{bmatrix}\omega_{1}&&&\\ &\omega_{2}&&\\ &&\ddots&\\ &&&\omega_{n}\end{bmatrix}\begin{bmatrix}\beta_{1}-\alpha_{1}\\ \beta_{2}-\alpha_{2}\\ \vdots\\ \beta_{n}-\alpha_{n}\end{bmatrix}.$
Since $\beta_{k}=(\Pi^{-1}\gamma)_{k}=\gamma_{k}+\gamma_{k+1}+\dots+\gamma_{n}>r$ and $\beta_{k+1}=\gamma_{k+1}+\gamma_{k+2}+\dots+\gamma_{n}\leq r$, we must have $\gamma_{k}>0$. By the KKT complementary slackness condition,
(9) $\lambda_{k}=2(\Pi^{-{\scriptscriptstyle\mathsf{T}}}W(\beta-\alpha))_{k}=2\sum_{i=1}^{k}\omega_{i}(\beta_{i}-\alpha_{i})=0.$
Hence we have
$\int(\operatorname{proj}_{G(\mathcal{R})}h)\cdot\mathbbm{1}_{E}\,\mathrm{d}\nu=\sum_{i=1}^{k}\beta_{i}\omega_{i}=\sum_{i=1}^{k}\alpha_{i}\omega_{i}=\int h\mathbbm{1}_{E}\,\mathrm{d}\nu.$
Step ii: Next we drop the ordered simple assumption and just require that $g\in L^{2}_{+}(\nu)$ be a general function with
(10) $\nu(\{\operatorname{proj}_{G(\mathcal{R})}g=r\})=0.$
Let $(h_{n})$ be a sequence of nonnegative ordered simple functions with $h_{n}\to g$ almost surely and in $L^{2}$-norm. So $\operatorname{proj}_{G(\mathcal{R})}h_{n}\to\operatorname{proj}_{G(\mathcal{R})}g$ in $L^{2}$-norm and thus in measure. By passing through a subsequence if necessary, we may assume that $\operatorname{proj}_{G(\mathcal{R})}h_{n}\to\operatorname{proj}_{G(\mathcal{R})}g$ almost surely and in $L^{2}$-norm. Set $E=\{\operatorname{proj}_{G(\mathcal{R})}g>r\}$ and $E_{n}=\{\operatorname{proj}_{G(\mathcal{R})}h_{n}>r\}$, bearing in mind that we assume (10). Then
$\lVert\mathbbm{1}_{E}-\mathbbm{1}_{E_{n}}\rVert^{2}$ $=\int(\mathbbm{1}_{E}-\mathbbm{1}_{E_{n}})^{2}\,\mathrm{d}\nu$
$=\nu(E)+\nu(E_{n})-2\nu(E\cap E_{n})=\nu(E\triangle E_{n})=\nu(E\cap E_{n}^{\mathsf{c}})+\nu(E_{n}\cap E^{\mathsf{c}}).$
Consider the null set $N=\{\operatorname{proj}_{G(\mathcal{R})}h_{n}\not\to\operatorname{proj}_{G(\mathcal{R})}g\}$. If
$x\in E\cap E_{n}^{\mathsf{c}}=\{\operatorname{proj}_{G(\mathcal{R})}g>r\}\cap\{\operatorname{proj}_{G(\mathcal{R})}h_{n}\leq r\}$
infinitely often, then
$\liminf_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}h_{n}(x)\leq r<\operatorname{proj}_{G(\mathcal{R})}g(x),$
and so $x\in N$. Consequently, $\limsup_{n\to\infty}E\cap E_{n}^{\mathsf{c}}\subseteq N$. On the other hand, we may decompose
$E_{n}\cap E^{\mathsf{c}}$ =[$=\bigl{[}{\{\operatorname{proj}_{G(\mathcal{R})}h_{n}>r\}\cap\{\operatorname{proj}_{G(\mathcal{R})}g<r\}}\bigr{]}\cup\bigl{[}\{\operatorname{proj}_{G(\mathcal{R})}h_{n}>r\}\cap\{\operatorname{proj}_{G(\mathcal{R})}g=r\}\bigr{]}$
$\eqqcolon F_{n}\cup G_{n}.$
Since $G_{n}$ is a null set for each $n$, if $x\in F_{n}$ infinitely often, then
$\limsup_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}h_{n}(x)\geq r>\operatorname{proj}_{G(\mathcal{R})}g(x),$
and so $x\in N$ as well. Consequently, $\limsup_{n\to\infty}F_{n}\subseteq N$. We obtain
$\limsup_{n\to\infty}\lVert\mathbbm{1}_{E}-\mathbbm{1}_{E_{n}}\rVert^{2}$ $\leq\limsup_{n\to\infty}\bigl{[}\nu(E\cap E_{n}^{\mathsf{c}})+\nu(F_{n})+\nu(G_{n})\bigr{]}$
$\leq\limsup_{n\to\infty}\nu(E\cap E_{n}^{\mathsf{c}})+\limsup_{n\to\infty}\nu(F_{n})$
$\leq\nu\Bigl{(}\limsup_{n\to\infty}E\cap E_{n}^{\mathsf{c}}\Bigr{)}+\nu\Bigl{(}\limsup_{n\to\infty}F_{n}\Bigr{)}=0,$
and by Hölder,
$\lVert h\mathbbm{1}_{E}-h_{n}\mathbbm{1}_{E_{n}}\rVert_{1}$ $=\lVert h\mathbbm{1}_{E}-h\mathbbm{1}_{E_{n}}+h\mathbbm{1}_{E_{n}}-h_{n}\mathbbm{1}_{E_{n}}\rVert_{1}$
$\leq\lVert h\cdot(\mathbbm{1}_{E}-\mathbbm{1}_{E_{n}})\rVert_{1}+\lVert(h-h_{n})\cdot\mathbbm{1}_{E_{n}}\rVert_{1}$
$\leq\lVert h\rVert\lVert\mathbbm{1}_{E}-\mathbbm{1}_{E_{n}}\rVert+\lVert h-h_{n}\rVert\lVert\mathbbm{1}_{E_{n}}\rVert\to 0.$
The same argument also yields
$(\operatorname{proj}_{G(\mathcal{R})}h_{n})\cdot\mathbbm{1}_{E_{n}}\to(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{E}$
in $L^{1}$-norm. By Lemma 2.2, we get
$\int g\mathbbm{1}_{E}\,\mathrm{d}\nu=\lim_{n\to\infty}\int h_{n}\mathbbm{1}_{E_{n}}\,\mathrm{d}\nu=\lim_{n\to\infty}\int(\operatorname{proj}_{G(\mathcal{R})}h_{n})\cdot\mathbbm{1}_{E_{n}}\,\mathrm{d}\nu=\int(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{E}\,\mathrm{d}\nu$
as required.
Step iii: We next omit the condition (10). Let $x=\inf\{\operatorname{proj}_{G(\mathcal{R})}g=r\}$ and assume $x>-\infty$ for nontriviality. For any $y<x$, we have $\operatorname{proj}_{G(\mathcal{R})}g(y)\geq r$ and so
(11) $\inf_{y\in(-\infty,x)}\operatorname{proj}_{G(\mathcal{R})}g(y)\geq r.$
If we have strict inequality in (11), then there exists $\varepsilon>0$ such that $\inf_{y\in(-\infty,x)}\operatorname{proj}_{G(\mathcal{R})}g(y)>r+\varepsilon$. Thus
$E_{r}\coloneqq\{\operatorname{proj}_{G(\mathcal{R})}g>r\}=\{\operatorname{proj}_{G(\mathcal{R})}g>r+\varepsilon\}\eqqcolon E_{r+\varepsilon}.$
Since $\operatorname{proj}_{G(\mathcal{R})}g\neq r+\varepsilon$ almost surely, $\nu(\{\operatorname{proj}_{G(\mathcal{R})}g=r+\varepsilon\})=0$, and
$\int_{E_{r}}\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu=\int_{E_{r+\varepsilon}}\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu=\int_{E_{r+\varepsilon}}g\,\mathrm{d}\nu=\int_{E_{r}}g\,\mathrm{d}\nu.$
If we have equality in (11), then for any $k\in\mathbb{N}$, we may pick $x_{k}\in(-\infty,x)$ such that
$r<\operatorname{proj}_{G(\mathcal{R})}g(x_{k})<r+\frac{1}{k}$
and $r_{k}\in(r,\operatorname{proj}_{G(\mathcal{R})}g(x_{k})]$ such that
$r<r_{k}<r+\frac{1}{k}\quad\text{and}\quad\nu(\operatorname{proj}_{G(\mathcal{R})}g=r_{k})=0.$
Hence $E_{k}\coloneqq\{\operatorname{proj}_{G(\mathcal{R})}g>r_{k}\}$, $k\in\mathbb{N}$, is an increasing sequence of nested sets with $E_{k}\subseteq E_{r}$ and so $\bigcup_{k=1}^{\infty}E_{k}\subseteq E_{r}$. Conversely, if $y\in E_{r}$, i.e., $\operatorname{proj}_{G(\mathcal{R})}g(y)>r$, then there exists $r_{k}$ such that $r_{k}<\operatorname{proj}_{G(\mathcal{R})}g(y)$; so $y\in E_{k}$; and so $E_{k}\uparrow E_{r}$. Define the finite measures
$\zeta(A)\coloneqq\int_{A}\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu,\qquad\xi(A)\coloneqq\int_{A}g\,\mathrm{d}\nu$
for any $A\in\mathcal{B}$. By the continuity of measure from below,
$\zeta(E_{r})=\lim_{k\to\infty}\zeta(E_{k})=\lim_{k\to\infty}\xi(E_{k})=\xi(E_{r})$
as required.
Step iv: It remains to treat the case $E=\mathbb{R}$. We will show that $\int\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu=\int g\,\mathrm{d}\nu$ by first establishing it for an ordered simple function $h=\sum_{i=1}^{n}\alpha_{i}\mathbbm{1}_{A_{i}}$. We may assume that $\beta_{k}>0$ and $\beta_{k+1}=\beta_{k+2}=\dots=\beta_{n}=0$ without loss of generality. By (8) and Step i,
$\lambda_{n}$ $=2\sum_{i=1}^{n}\omega_{i}(\beta_{i}-\alpha_{i})=2\sum_{i=1}^{k}\omega_{i}(\beta_{i}-\alpha_{i})+2\sum_{j=k+1}^{n}\omega_{i}(\beta_{i}-\alpha_{i})$
$=-\lambda_{k}-2(\omega_{k+1}\alpha_{k+1}+\omega_{k+2}\alpha_{k+2}+\dots+\omega_{n}\alpha_{n}).$
By (9), $\lambda_{k}=0$. By KKT dual feasibility, $\lambda_{n}\geq 0$. Since we also have $\omega_{i},\alpha_{i}\geq 0$, $i=k+1,k+2,\dots,n$,
$\lambda_{n}=-2(\omega_{k+1}\alpha_{k+1}+\omega_{k+2}\alpha_{k+2}+\dots+\omega_{n}\alpha_{n})=0.$
It follows from (8) that $0=\lambda_{n}=2\sum_{i=1}^{n}\omega_{i}(\beta_{i}-\alpha_{i})$ and so
$\int\operatorname{proj}_{G(\mathcal{R})}h\,\mathrm{d}\nu=\sum_{i=1}^{n}\beta_{i}w_{i}=\sum_{i=1}^{n}\alpha_{i}w_{i}=\int h\,\mathrm{d}\nu.$
For a general function $g\in L^{2}_{+}(\nu)$, pick a sequence of nonnegative ordered simple functions $h_{n}\to g$ in $L^{2}$-norm and apply Lemma 2.1 to get
$\int g\,\mathrm{d}\nu=\lim_{n\to\infty}\int h_{n}\,\mathrm{d}\nu=\lim_{n\to\infty}\int\operatorname{proj}_{G(\mathcal{R})}h_{n}\,\mathrm{d}\nu=\int\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu.\qed$
We state an immediate corollary of Theorem 3.9.
Corollary 3.10.
Let $g\in L^{2}_{+}(\nu)$. If $\operatorname{proj}_{G(\mathcal{R})}g(E)\in\mathcal{B}$, then
$\int_{E}\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu=\int_{E}g\,\mathrm{d}\nu.$
Proof.
Theorem 3.9 says that (7) holds for any $E\in\mathcal{E}(g)$. The left and right sides of (7) define finite measures $\zeta$ and $\xi$ respectively, and $\zeta(\mathbb{R})=\xi(\mathbb{R})$ as in Step iii of the proof of Theorem 3.9. Since $\mathcal{E}(g)$ is a $\pi$-system, $\sigma(\mathcal{E}(g))=(\operatorname{proj}_{G(\mathcal{R})}g)^{-1}(\mathcal{B})$ by Lemmas 2.3 and 2.4. Hence (7) holds for all $E\in(\operatorname{proj}_{G(\mathcal{R})}g)^{-1}(\mathcal{B})$. ∎
We now arrive at the Radon–Nikodym property for $\mathcal{R}$ mentioned earlier.
Corollary 3.11 (Radon–Nikodym over rays).
Let $\mu,\nu$ be Borel probability measures on $(\mathbb{R},\mathcal{B})$. Then the $\mathcal{R}$-measurable function
$\rho_{\mathcal{R}}=\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu},$
and the $\pi$-system $\mathcal{E}(\mathrm{d}\mu/\mathrm{d}\nu)=\bigl{\{}\{\rho_{\mathcal{R}}>r\}:r\geq 0\bigr{\}}\cup\{\mathbb{R}\}\subseteq\mathcal{R}$ satisfy
$\mu(E)=\int_{E}\rho_{\mathcal{R}}\,\mathrm{d}\nu$
for any $E\in\mathcal{E}(\mathrm{d}\mu/\mathrm{d}\nu)$.
Proof.
For any $E\in\mathcal{E}(\mathrm{d}\mu/\mathrm{d}\nu)$, it follows from Theorem 3.9 that
$\mu(E)=\int_{E}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu=\int_{E}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu=\int_{E}\rho_{\mathcal{R}}\,\mathrm{d}\nu.\qed$
We will show that the following definition of $f$-divergence over $\mathcal{R}$ provides an affirmative answer to Question 1: Theorem 3.15 shows that (a) when $f$ is chosen to be $f(t)=\lvert t-1\rvert/2$, we recover the Kolmogorov–Smirnov distance in (3); and Definition 3.12 shows that by replacing $\mathcal{R}$ by $\mathcal{B}$, we recover the standard $f$-divergence in (5). A reminder of our convention in Section 2.3: we assume that $\mu\ll\nu$ whenever we write $\mathrm{d}\mu/\mathrm{d}\nu$. This is not an additional imposition; it is also a requirement for standard $f$-divergence in Definition 1.5.
Definition 3.12 ($f$-divergence over rays).
Let $\mu$ and $\nu$ be Borel probability measures and $\mathcal{R}$ be the class of rays. Then the $f$-divergence over $\mathcal{R}$ is defined as
(12) $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)\coloneqq\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu.$
Henceforth we will use this terminology in conjunction with standard nomenclatures. For example when we speak of total variation or Hellinger distance over $\mathcal{R}$ we mean the $f$-divergence over $\mathcal{R}$ with $f(t)=\lvert t-1\rvert/2$ or $f(t)=(\sqrt{t}-1)^{2}$ respectively. Obviously, if we replace $\mathcal{R}$ in (12) by the Borel $\sigma$-algebra $\mathcal{B}$, then $\operatorname{D}_{f}^{\mathcal{B}}(\mu\|\nu)=\operatorname{D}_{f}(\mu\|\nu)$ and we recover the standard $f$-divergence in Definition 1.5. So we may view the standard definition as “$f$-divergence over $\mathcal{B}$.”
We will provide abundant evidence in Section 4 that our $f$-divergence over $\mathcal{R}$ in Definition 3.12 retains almost all known properties of the standard $f$-divergence. But our most rudimentary justification is Theorem 3.15, i.e., the total variation distance on $\mathcal{R}$ is exactly the Kolmogorov–Smirnov distance. To get there we need two corollaries.
Corollary 3.13.
Let $g\in L^{2}_{+}(\nu)$. Then for any $A\in\mathcal{R}$,
$\int_{A}\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu\geq\int_{A}g\,\mathrm{d}\nu.$
Proof.
By Proposition 3.8 and Theorem 3.9, we have
$\int_{A}\operatorname{proj}_{G(\mathcal{R})}g\,\mathrm{d}\nu=\int(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{A}\,\mathrm{d}\nu\geq\int\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{A})\,\mathrm{d}\nu=\int_{A}g\,\mathrm{d}\nu.\qed$
The $\pi$-system $\mathcal{E}(g)$ as defined in (6) has the following property.
Corollary 3.14.
Let $g\in L^{2}_{+}(\nu)$. Then for any $E\in\mathcal{E}(g)$,
$(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{E}=\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{E}).$
Proof.
By Theorem 3.9,
$\int(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{E}\,\mathrm{d}\nu=\int g\mathbbm{1}_{E}\,\mathrm{d}\nu=\int\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{E})\,\mathrm{d}\nu.$
Since $(\operatorname{proj}_{G(\mathcal{R})}g)\cdot\mathbbm{1}_{E}\geq\operatorname{proj}_{G(\mathcal{R})}(g\mathbbm{1}_{E})$, we must in fact have equality. ∎
We now arrive at the main result of this section. The key to Theorem 3.15 is Theorem 3.9. If we have an analogue of Theorem 3.9 for a class of sets $\mathcal{C}$, then we have an “$f$-divergence on $\mathcal{C}$” for which an analogue of Theorem 3.15 with $\mathcal{C}$ in place of $\mathcal{R}$ holds.
Theorem 3.15 (Kolmogorov–Smirnov as special case).
Let $\mu,\nu$ be Borel probability measures on $(\mathbb{R},\mathcal{B})$. Then
$\int\frac{1}{2}\Bigl{\lvert}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{\rvert}\,\mathrm{d}\nu=\sup_{A\in\mathcal{R}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}.$
Proof.
Let $E=\{\operatorname{proj}_{G(\mathcal{R})}\mathrm{d}\mu/\mathrm{d}\nu>1\}$. By Theorem 3.9 and Corollary 3.11,
$\int_{A}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu=\int_{A}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu=\mu(A)$
for any $A\in\mathcal{E}(\mathrm{d}\mu/\mathrm{d}\nu)$. So
$\int\frac{1}{2}\Bigl{\lvert}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{\rvert}\,\mathrm{d}\nu$ $=\frac{1}{2}\int_{E}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\,\mathrm{d}\nu+\frac{1}{2}\int_{E^{\mathsf{c}}}1-\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu$
$=\frac{1}{2}\Bigl{[}2\int_{E}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu-1\Bigr{]}-\frac{1}{2}\Bigl{[}2\int_{E}\,\mathrm{d}\nu-1\Bigr{]}$
(13) $=\int_{E}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu-\nu(E)$
$=\mu(E)-\nu(E)\leq\sup_{A\in\mathcal{R}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)},$
noting that $E=\{\operatorname{proj}_{G(\mathcal{R})}\mathrm{d}\mu/\mathrm{d}\nu>1\}\in\mathcal{R}$ as $\operatorname{proj}_{G(\mathcal{R})}\mathrm{d}\mu/\mathrm{d}\nu\in G(\mathcal{R})$. For the reverse inequality, by Corollary 3.13, we have
$\int_{A}\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{)}\mathrm{d}\nu\geq\mu(A)-\nu(A),$
for any $A\in\mathcal{R}$. By (13),
$\int\frac{1}{2}\Bigl{\lvert}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{\rvert}\,\mathrm{d}\nu$ $=\int_{E}\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{)}\mathrm{d}\nu$
$=\sup_{A\in\mathcal{R}}\Bigl{(}\int_{A}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu-\nu(A)\Bigr{)}\geq\sup_{A\in\mathcal{R}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}$
as required. ∎
4. Properties of the $f$-divergences over rays
Theorem 3.15 serves as an indication that our notion of $f$-divergence over rays in Definition 3.12 is the right one because it restricts to the Kolmogorov–Smirnov distance for a special choice of $f$. In this section, we will see that just about every property that we know holds for $f$-divergences also holds for $f$-divergences over $\mathcal{R}$. Henceforth we will denote the convex cone of functions convex on $[0,\infty)$ and both vanishing and strictly convex at $1$ by
(14) $\mathcal{F}\coloneqq\{f:[0,\infty)\to\mathbb{R}:f\text{ is convex, }f\text{ isstrictly convex at 1, and }f(1)=0\}.$
The strict convexity at $1$ is as in [40, p. 120]: For all $x,y\in[0,\infty)$ and $t\in(0,1)$ such that $tx+(1-t)y=1$, we have $tf(x)+(1-t)f(y)>0$. This condition guarantees that $\operatorname{D}_{f}(\mu\|\nu)=0$ if and only if $\mu=\nu$. As a precaution, in situations like the Kullback–Liebler divergence where $f(x)=x\log x$, the value $f(0)$ may only be defined in a limiting sense.
Again $\mu,\nu$ will denote Borel probability measures on $(\mathbb{R},\mathcal{B})$ throughout this section, with the implicit assumption that $\mu\ll\nu$ whenever we speak of their Radon–Nikodym derivative $\mathrm{d}\mu/\mathrm{d}\nu$.
Theorem 4.1.
Let $f,g\in\mathcal{F}$. Then the $f$-divergence over $\mathcal{R}$ has the following properties:
[…]
    Linearity: $\operatorname{D}_{\alpha f+\beta g}^{\mathcal{R}}(\mu\|\nu)=\alpha\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)+\beta\operatorname{D}_{g}^{\mathcal{R}}(\mu\|\nu)$ for any $\alpha,\beta\geq 0$.
[…]
    Nonnegativity: $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)\geq 0$.
[…]
    Affine invariance: If $g(x)=f(x)+c(x-1)$, $c\in\mathbb{R}$, then $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)=\operatorname{D}_{g}^{\mathcal{R}}(\mu\|\nu)$.
[…]
    Boundedness: $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)\leq\operatorname{D}_{f}(\mu\|\nu)$.
[…]
    Identity: If $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)=0=\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\mu)$, then $\mu=\nu$.
Proof.
Item i is routine from definition. For item ii, since $f$ is convex, applying Jensen’s inequality and Theorem 3.9 gives
$\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu\geq f\Bigl{(}\int\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu\Bigr{)}=f\Bigl{(}\int\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu\Bigr{)}=f(1)=0.$
Item iii follows from
$\operatorname{D}_{g}^{\mathcal{R}}(\mu\|\nu)$ $=\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}+c\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{)}\,\mathrm{d}\nu$
$=\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu+c\int\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{)}\mathrm{d}\nu=\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu),$
noting that the last integral vanishes. The proofs for items iv and v are considerably more involved and will be deferred to Proposition 4.3 and Corollary 4.7 respectively. ∎
We caution the reader that special properties of a specific $f$-divergence may sometimes be lost for its counterpart over $\mathcal{R}$. For example, it is well-known that while $f$-divergence is generally not symmetric, i.e., $\operatorname{D}_{f}(\mu\|\nu)\neq\operatorname{D}_{f}(\nu\|\mu)$, the total variation distance is. This symmetry is lost for the total variation distance over $\mathcal{R}$ as
$\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\mu\|\nu)=\int\frac{1}{2}\Bigl{\lvert}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\Bigr{\rvert}\,\mathrm{d}\nu=\sup_{A\in\mathcal{R}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)};$
although one may recover it by symmetrizing:
$\widehat{\operatorname{D}}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\mu\|\nu)\coloneqq\max\bigl{\{}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\mu\|\nu),\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\nu\|\mu)\bigr{\}}=\sup_{A\in\mathcal{R}}\,\lvert\mu(A)-\nu(A)\rvert.$
This is also the reason why we need both $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)=0$ and $\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\mu)=0$ in Theorem 4.1v. It would not be true otherwise. We illustrate this in Figure 1 with the total variation and Hellinger distances over $\mathcal{R}$. One can see that $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)=0$ or $\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\mu)=0$ does not necessarily imply $\mu=\nu$ but that $\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)=\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\mu)=0$ does.
(a) $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)$
(b) $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\mu\|\nu)$
(c) $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\nu\|\mu)$
(d) $\operatorname{D}_{\operatorname{\scriptscriptstyle H}}(\mu\|\nu)$
(e) $\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{\mathcal{R}}(\mu\|\nu)$
(f) $\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{\mathcal{R}}(\nu\|\mu)$
Figure 1. 40 Level curves of total variation $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}$, total variation over rays $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}$, Hellinger distance $\operatorname{D}_{\operatorname{\scriptscriptstyle H}}$ and Hellinger distance over rays $\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{\mathcal{R}}$ for fixed $\nu=[0.2,0.5,0.3]$ as $\mu$ ranges over the simplex of distributions on a three-element set.
Evidently, the projection operator $\operatorname{proj}_{G(\mathcal{R})}$ plays a key role in our definition of $f$-divergence over $\mathcal{R}$. The next two results show how it interacts with nondecreasing functions and with convex functions.
Lemma 4.2.
Let $f:\mathbb{R}\to\mathbb{R}$ be nondecreasing. If $g\in L^{2}_{+}(\nu)$ is almost surely continuous, then
$\int(\operatorname{proj}_{G(\mathcal{R})}g)\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}g)\,\mathrm{d}\nu=\int g\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}g)\,\mathrm{d}\nu.$
Proof.
Let $h=\sum_{i=1}^{n}\alpha_{i}\mathbbm{1}_{A_{i}}$ and $\operatorname{proj}_{G(\mathcal{R})}h=\sum_{i=1}^{n}\beta_{i}\mathbbm{1}_{A_{i}}$. Suppose $\operatorname{proj}_{G(\mathcal{R})}h$ takes $m$ distinct values $\tau_{1}>\dots>\tau_{m}$ where $m\leq n$. For each $k=1,\dots,m$, set
$E_{k}\coloneqq\bigcup_{i=1}^{k}\{\operatorname{proj}_{G(\mathcal{R})}h=\tau_{i}\}\in\mathcal{E}(h).$
Let $\zeta_{m}\coloneqq f(\tau_{m})$ and $\zeta_{m-j}\coloneqq f(\tau_{m-j})-\zeta_{m-j+1}-\zeta_{m-j+2}-\dots-\zeta_{m}$, $j=1,\dots,m-1$. Then
$\int(\operatorname{proj}_{G(\mathcal{R})}h)\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}h)\,\mathrm{d}\nu$ $=\int(\operatorname{proj}_{G(\mathcal{R})}h)\cdot\Bigl{(}\sum_{i=1}^{n}f(\beta_{i})\mathbbm{1}_{A_{i}}\Bigr{)}\,\mathrm{d}\nu$
$=\int(\operatorname{proj}_{G(\mathcal{R})}h)\cdot\sum_{i=1}^{m}\zeta_{i}\mathbbm{1}_{E_{i}}\,\mathrm{d}\nu$
$=\sum_{i=1}^{m}\zeta_{i}\int_{E_{i}}\operatorname{proj}_{G(\mathcal{R})}h\,\mathrm{d}\nu=\sum_{i=1}^{m}\zeta_{i}\int_{E_{i}}h\,\mathrm{d}\nu$
$=\int h\cdot\sum_{i=1}^{m}\zeta_{i}\mathbbm{1}_{E_{i}}\,\mathrm{d}\nu=\int h\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}h)\,\mathrm{d}\nu.$
Now let $h_{n}$ be a nondecreasing sequence of nonnegative ordered simple function that converges to $g$ almost surely and in $L^{2}$-norm. By passing through a subsequence if necessary, $\operatorname{proj}_{G(\mathcal{R})}h_{n}$ is a nonincreasing sequence that converges to $\operatorname{proj}_{G(\mathcal{R})}g$ almost surely. Since $f$ is nondecreasing,
$f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}h_{n}\uparrow f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}g$
almost surely. By the monotone convergence theorem,
$\int(\operatorname{proj}_{G(\mathcal{R})}g)\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}g)\,\mathrm{d}\nu=\lim_{n\to\infty}\int(\operatorname{proj}_{G(\mathcal{R})}h_{n})\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}h_{n})\,\mathrm{d}\nu\\ =\lim_{n\to\infty}\int h_{n}\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}h_{n})\,\mathrm{d}\nu=\int g\cdot(f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}\operatorname{proj}_{G(\mathcal{R})}g)\,\mathrm{d}\nu.\qed$
Proposition 4.3.
Let $f:\mathbb{R}\to\mathbb{R}$ be convex. If $g\in L^{2}_{+}(\nu)$ is almost surely continuous, then
$\int f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}{\operatorname{proj}_{G(\mathcal{R})}g}\,\mathrm{d}\nu\leq\int f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}g\,\mathrm{d}\nu.$
Proof.
By convexity,
$f\bigl{(}\operatorname{proj}_{G(\mathcal{R})}g(x)\bigr{)}-f\bigl{(}g(x)\bigr{)}\leq f^{\prime}\bigl{(}\operatorname{proj}_{G(\mathcal{R})}g(x)\bigr{)}\cdot\bigl{(}\operatorname{proj}_{G(\mathcal{R})}g(x)-g(x)\bigr{)}$
almost surely. Integrating with respect to $\nu$ and applying Lemma 4.2 to $f^{\prime}$, we get
$\int f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}{\operatorname{proj}_{G(\mathcal{R})}g}-f\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}g\,\mathrm{d}\nu\leq\int f^{\prime}(\operatorname{proj}_{G(\mathcal{R})}g)\cdot(\operatorname{proj}_{G(\mathcal{R})}g-g)\,\mathrm{d}\nu=0.\qed$
Theorem 4.1iv follows from Proposition 4.3 since any $f\in\mathcal{F}$ is convex and so
$\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\mu)=\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu\leq\int f\Bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu=\operatorname{D}_{f}(\nu\|\mu).$
It remains to establish Theorem 4.1v, whose proof will take four intermediate results: Propositions 4.4 and 4.5, Corollaries 4.6 and 4.7, each of independent interest. We begin by showing how we may convert $f$-divergence over rays to standard $f$-divergence.
Proposition 4.4.
For any Borel probability measures $\mu$ and $\nu$, there exists a Borel probability measure $\zeta$ such that for any $f\in\mathcal{F}$,
$\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)=\operatorname{D}_{f}(\zeta\|\nu).$
Proof.
We define $\zeta$ by
$\zeta(B)\coloneqq\int_{B}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu$
for any $B\in\mathcal{B}$. By Corollary 3.11, $\zeta$ is a probability measure absolutely continuous with respect to $\nu$. Hence
$\frac{\mathrm{d}\zeta}{\mathrm{d}\nu}=\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$
and so
$\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)=\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu=\int f\Bigl{(}\frac{\mathrm{d}\zeta}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu=\operatorname{D}_{f}(\zeta\|\nu).\qed$
Next we do the reverse: converting standard $f$-divergence to $f$-divergence over rays.
Proposition 4.5.
For any Borel probability measures $\mu$ and $\nu$, there exist Borel probability measures $\eta,\tau$ such that for any $f\in\mathcal{F}$,
$\operatorname{D}_{f}(\mu\|\nu)=\operatorname{D}_{f}^{\mathcal{R}}(\eta\|\tau).$
Proof.
The distribution function $m:[0,\infty)\to[0,1]$ of $\mathrm{d}\mu/\mathrm{d}\nu$ is given by
$m(\alpha)\coloneqq\nu\Bigl{(}\Bigl{\{}x\in X:\frac{\mathrm{d}\mu}{\mathrm{d}\nu}(x)>\alpha\Bigr{\}}\Bigr{)}$
for any $\alpha\in[0,\infty)$. Since $m$ may not be injective, we define its generalized inverse $\rho:[0,1]\to[0,\infty)$,
$\rho(u)\coloneqq\inf\{\alpha\geq 0:m(\alpha)\leq u\}$
for any $u\in[0,1]$. Thus $\rho$ may be viewed a decreasing rearrangement of $\mathrm{d}\mu/\mathrm{d}\nu$ on $[0,1]$ satisfying
$\nu\Bigl{(}\Bigl{\{}x:\frac{\mathrm{d}\mu}{\mathrm{d}\nu}(x)>\alpha\Bigr{\}}\Bigr{)}=\lambda\bigl{(}\{u:\rho(u)>\alpha\}\bigr{)},$
with $\lambda$ the Lebesgue measure on $[0,1]$ [20, p. 199]. Let $T:[0,1]\to\mathbb{R}$ be a strictly increasing measure-preserving transformation. Define the probability measures
$\tau(A)\coloneqq\lambda(T^{-1}(A)),\qquad\eta(A)\coloneqq\int_{T^{-1}(A)}\rho\,\mathrm{d}\lambda$
for any $A\in\mathcal{B}$. Since $\rho$ is nonincreasing and $T^{-1}$ is increasing, the Radon–Nikodym derivative
$\frac{\mathrm{d}\eta}{\mathrm{d}\tau}=\rho\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}T^{-1}$
is also nonincreasing on $\mathbb{R}$, so $\mathrm{d}\eta/\mathrm{d}\tau=\operatorname{proj}_{G(\mathcal{R})}\mathrm{d}\eta/\mathrm{d}\tau$. Hence, for any $f\in\mathcal{F}$,
$\operatorname{D}_{f}(\mu\|\nu)=\int f\Bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu=\int_{[0,1]}f(\rho)\,\mathrm{d}\lambda=\int f(\rho\mathbin{\mathchoice{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}{\vbox{\hbox{$\scriptscriptstyle\circ$}}}}T^{-1})\,\mathrm{d}\tau=\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\eta}{\mathrm{d}\tau}\Bigr{)}\,\mathrm{d}\tau=\operatorname{D}_{f}^{\mathcal{R}}(\eta\|\tau)$
as required. ∎
Collectively, Propositions 4.4 and 4.5 lead to the following conclusion: An inequality relation between two divergences hold if and only if that same relation hold for their counterparts over rays.
Corollary 4.6 (Preservation of inequality relations).
Let $f,g\in\mathcal{F}$ and $\psi,\varphi:[0,\infty)\to\mathbb{R}$. Then
$\psi(\operatorname{D}_{f}(\mu\|\nu))\leq\varphi(\operatorname{D}_{g}(\mu\|\nu))$
holds for all Borel probability measures $\mu$ and $\nu$ if and only if
$\psi(\operatorname{D}^{\mathcal{R}}_{f}(\mu\|\nu))\leq\varphi(\operatorname{D}^{\mathcal{R}}_{g}(\mu\|\nu))$
holds for all Borel probability measures $\mu$ and $\nu$.
Proof.
Suppose $\psi(\operatorname{D}_{f}(\mu\|\nu))\leq\varphi(\operatorname{D}_{g}(\mu\|\nu))$ for all $\mu$ and $\nu$. By Proposition 4.4, there exists $\zeta$ such that
$\psi(\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu))=\psi(\operatorname{D}_{f}(\zeta\|\nu))\leq\varphi(\operatorname{D}_{g}(\zeta\|\nu))=\varphi(\operatorname{D}_{g}^{\mathcal{R}}(\mu\|\nu)).$
Conversely, suppose $\psi(\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu))\leq\varphi(\operatorname{D}^{\mathcal{R}}_{g}(\mu\|\nu))$ for all $\mu$ and $\nu$. By Proposition 4.5, there exist $\eta$ and $\tau$ such that
$\psi(\operatorname{D}_{f}(\mu\|\nu))=\psi(\operatorname{D}_{f}^{\mathcal{R}}(\eta\|\tau))\leq\varphi(\operatorname{D}_{g}^{\mathcal{R}}(\eta\|\tau))=\varphi(\operatorname{D}_{g}(\mu\|\nu)).\qed$
As most readers familiar with $f$-divergences would suspect, the total variation distance plays a somewhat special role among all $f$-divergences. One striking property [7] is that it serves as a kind of universal lower bound for all $f$-divergence, in the sense that
$\bar{f}\bigl{(}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)\bigr{)}\leq\operatorname{D}_{f}(\mu\|\nu),$
where $\bar{f}(x)\coloneqq f(1+x)+f(1-x)$ is the symmetrized function of $f$. Note that $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)$ takes value in $[0,1]$, so the $f(1-x)$ term is always evaluated within the domain of $f\in\mathcal{F}$. A consequence of our definition of $\mathcal{F}$ in (14) is that $\bar{f}$ is increasing and $\bar{f}(x)>0$ when $x>0$. So by Corollary 4.6, the same result holds verbatim for divergences over rays.
Corollary 4.7 (Total variation over rays as universal lower bound).
Let $f\in\mathcal{F}$ and set $\bar{f}(x)\coloneqq f(1+x)+f(1-x)$ for $x\in[0,1]$. Then
$\bar{f}\bigl{(}\operatorname{D}^{\mathcal{R}}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)\bigr{)}\leq\operatorname{D}^{\mathcal{R}}_{f}(\mu\|\nu).$
Theorem 4.1v now follows directly from Corollary 4.7: If $\operatorname{D}^{\mathcal{R}}_{f}(\mu\|\nu)=0$, then $\bar{f}\bigl{(}\operatorname{D}^{\mathcal{R}}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)\bigr{)}=0$. So $\operatorname{D}^{\mathcal{R}}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)=0$ as $\bar{f}$ is strictly increasing and $\bar{f}(0)=0$. Since we have also assumed that $\operatorname{D}^{\mathcal{R}}_{f}(\nu\|\mu)=0$, we get $\operatorname{D}^{\mathcal{R}}_{\operatorname{\scriptscriptstyle TV}}(\nu\|\mu)=0$, and thus
$\sup_{A\in\mathcal{R}}\,\lvert\mu(A)-\nu(A)\rvert=\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\mu\|\nu)\vee\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\nu\|\mu)=0.$
This shows that $\mu=\nu$ over $\mathcal{R}$. But $\mathcal{R}$ is a $\pi$-system that generates $\mathcal{B}$. So Lemma 2.3 ensures that we actually have $\mu=\nu$.
We end this section with another demonstration of the utility of Corollary 4.6: It allows us to extend essentially all known relations between divergences to their counterparts over rays.
Theorem 4.8 (Relations between $f$-divergences over rays).
The following inequalities hold verbatim when each $\operatorname{D}_{f}$ is replaced by $\operatorname{D}_{f}^{\mathcal{R}}$.
[…]
    Total variation and Hellinger:
[…]
    Kullback–Leibler and total variation:
    $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{2}(\mu\|\nu)\leq\frac{1}{2}\operatorname{D}_{\operatorname{\scriptscriptstyle KL}}(\mu\|\nu),$
    $\log\Bigl{(}\frac{1+\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)}{1-\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)}\Bigr{)}-\frac{2\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)}{1+\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)}\leq\operatorname{D}_{\operatorname{\scriptscriptstyle KL}}(\mu\|\nu),$
    $\operatorname{D}_{\operatorname{\scriptscriptstyle KL}}(\mu\|\nu)\leq\log\Bigl{(}1+\frac{1}{2\nu_{\min}}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)^{2}\Bigr{)}\leq\frac{1}{2\nu_{\min}}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)^{2},$
    where $\nu_{\min}\coloneqq\min_{x}\nu(x)$. These are known respectively as Pinsker, strong Pinsker, and reverse Pinsker inequality.
[…]
    Total variation and $\chi^{2}$:
    $\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{2}(\mu\|\nu)\leq\frac{1}{4}\operatorname{D}_{\chi^{2}}(\mu,\nu),\qquad\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu)\leq\max\biggl{\{}\frac{1}{2},\frac{\operatorname{D}_{\chi^{2}}(\mu\|\nu)}{1+\operatorname{D}_{\chi^{2}}(\mu,\nu)}\biggr{\}},$
    $4\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{2}(\mu\|\nu)\leq\varphi(\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\mu\|\nu))\leq\operatorname{D}_{\chi^{2}}(\mu\|\nu),$
    where
    $\varphi(t)=\begin{cases}4t^{2}&t\leq 1/2,\\ t/(1-t)&t\geq 1/2.\end{cases}$
[…]
    Kullback–Leibler and Hellinger:
    $\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{2}(\mu\|\nu)\leq 2\log\frac{2}{2-\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{2}(\mu\|\nu)}\leq\operatorname{D}_{\operatorname{\scriptscriptstyle KL}}(\mu\|\nu),$
    $\operatorname{D}_{\operatorname{\scriptscriptstyle KL}}(\mu\|\nu)\leq\frac{\log\bigl{(}1/\nu_{\min}-1\bigr{)}}{1-2\nu_{\min}}\bigl{(}1-(1-\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{2}(\mu\|\nu))^{2}\bigr{)}.$
[…]
    Kullback–Leibler and $\chi^{2}$:
    $\operatorname{D}_{\operatorname{\scriptscriptstyle KL}}(\mu\|\nu)\leq\log\bigl{(}1+\operatorname{D}_{\chi^{2}}(\mu\|\nu)\bigr{)}\leq\operatorname{D}_{\chi^{2}}(\mu\|\nu).$
[…]
    Le Cam and Hellinger:
    $\frac{1}{2}\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{2}(\mu\|\nu)\leq\operatorname{D}_{\operatorname{\scriptscriptstyle LC}}(\mu\|\nu)\leq\operatorname{D}_{\operatorname{\scriptscriptstyle H}}^{2}(\mu\|\nu).$
[…]
    Le Cam and Jensen–Shannon:
    $\operatorname{D}_{\operatorname{\scriptscriptstyle LC}}(\mu\|\nu)\leq\operatorname{D}_{\operatorname{\scriptscriptstyle JS}}(\mu\|\nu)\leq 2\log 2\cdot\operatorname{D}_{\operatorname{\scriptscriptstyle LC}}(\mu\|\nu).$
Proof.
These are well-known inequalities between the respective $f$-divergences: the two inequalities in i may be found in [40, p. 124] and [21, Theorem 12] respectively; the three inequalities in ii in [52, p. 88], [53], and [42, Theorem 28] respectively; those in iii, iv, and v may be found in [40, p. 133]; vi in [32, p. 48]; vii in [50, Theorem 3.2]. It follows from Corollary 4.6 that these relations also hold for the corresponding $f$-divergences over $\mathcal{R}$. ∎
Ultimately, a definition is justified by the richness of results that can be proved with it. We view the results in this section as evidence that the $f$-divergence over rays as defined in Definition 3.12 is the “right” one. In this regard, the next section will supply another major piece of evidence.
5. Glivenko–Cantelli theorem for $f$-divergences over rays
Put in our context, the original Glivenko–Cantelli theorem in Theorem 1.4 states that empirical distributions converge to the target distribution almost surely in total variation distance over $\mathcal{R}$, as defined in Definition 3.12. Here we generalize this result to all $f$-divergences over $\mathcal{R}$.
As in the case of total variation distance, this convergence does not always hold if we replace $\mathcal{R}$ with $\mathcal{B}$ or some other subsets of $\mathcal{B}$. Examples 1.1 and 1.2, where $\limsup_{n\to\infty}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\nu_{n}\|\nu)>0$, can be readily extended to any $f\in\mathcal{F}$ by way of Corollary 4.7:
$0<\bar{f}\bigl{(}\limsup_{n\to\infty}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\nu_{n}\|\nu)\bigr{)}=\limsup_{n\to\infty}\bar{f}\bigl{(}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\nu_{n}\|\nu)\bigr{)}\leq\operatorname{D}_{f}(\nu_{n}\|\nu),$
i.e., the $f$-divergence between the empirical distribution and the target does not in general converge to 00 almost surely. Our main result is that it holds for $f$-divergence over rays: For any $f\in\mathcal{F}$, Theorems 5.2 and 5.4 collectively show that almost surely
(15) $\lim_{n\to\infty}\operatorname{D}_{f}^{\mathcal{R}}(\nu_{n}\|\nu)=0=\lim_{n\to\infty}\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\nu_{n}).$
Note that an $f$-divergence is generally not symmetric in its arguments so we need to establish both equalities.
We will examine more carefully why a sequence of empirical measures can converge on $\mathcal{R}$ but fails to converge on $\mathcal{B}$. It turns out that a necessary condition is that its corresponding sequence of Radon–Nikodym derivatives does not converge to $1$, but the sequence of their projections onto $G(\mathcal{R})$ does.
Proposition 5.1.
Let $\nu$ be a Borel probability measure and $\nu_{n}$ be the corresponding empirical measure, $n\in\mathbb{N}$. Suppose the Radon–Nikodym derivative $\mathrm{d}\nu_{n}/\mathrm{d}\nu$ exists for all $n\in\mathbb{N}$ and are uniformly bounded. If
$\sup_{A\in\mathcal{B}}\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}\not\to 0\quad\text{and}\quad\sup_{A\in\mathcal{R}}\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}\to 0,$
then
$\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\not\to 1\quad\text{and}\quad\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\to 1.$
Proof.
Suppose $\lim_{n\to\infty}\mathrm{d}\nu_{n}/\mathrm{d}\nu=1$. Then
$0=\int\frac{1}{2}\biggl{\lvert}\lim_{n\to\infty}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}-1\biggr{\rvert}\,\mathrm{d}\nu\geq\limsup_{n\to\infty}\int\frac{1}{2}\biggl{\lvert}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}-1\biggr{\rvert}\,\mathrm{d}\nu=\limsup_{n\to\infty}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\nu_{n}\|\nu)\geq 0$
So $\lim_{n\to\infty}\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\nu_{n}\|\nu)=0$, a contradiction, and thus $\mathrm{d}\nu_{n}/\mathrm{d}\nu\not\to 1$. By assumption, $\mathrm{d}\nu_{n}/\mathrm{d}\nu\leq c$ for all $n\in\mathbb{N}$ and for some constant $c>0$. It then follows from Proposition 3.6 that
$\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\leq\operatorname{proj}_{G(\mathcal{R})}c=c$
for all $n$, i.e., the sequence $(\operatorname{proj}_{G(\mathcal{R})}\mathrm{d}\nu_{n}/\mathrm{d}\nu)_{n}$ is also uniformly bounded. Since $\dim_{\operatorname{\scriptscriptstyle VC}}(\mathcal{R})<\infty$, it follows from the Vapnik–Chervonenkis theorem (see p. 1 for a statement) that
$\lim_{n\to\infty}\operatorname{D}^{\mathcal{R}}_{\operatorname{\scriptscriptstyle TV}}(\nu_{n}\|\nu)=\lim_{n\to\infty}\sup_{A\in\mathcal{R}}\,\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}=0$
almost surely; taken together with the bounded convergence theorem,
00 0 $=\lim_{n\to\infty}\sup_{A\in\mathcal{R}}\,\bigl{(}\nu_{n}(A)-\nu(A)\bigr{)}=\lim_{n\to\infty}\int\frac{1}{2}\Bigl{\lvert}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}-1\Bigr{\rvert}\,\mathrm{d}\nu$
$=\int\frac{1}{2}\Bigl{\lvert}\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}-1\Bigr{\rvert}\,\mathrm{d}\nu\geq 0.$
Hence the last integrand must converge to zero, or, equivalently,
(16) $\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}=1.\qed$
With the observation in Proposition 5.1, we obtain the left half of (15).
Theorem 5.2 (Glivenko–Cantelli theorem for $f$-divergence I).
Let $\mathcal{R}$ be the class of rays, $f\in\mathcal{F}$ as in (14), $\nu$ a Borel probability measure, and $\nu_{n}$ the corresponding empirical measure, $n\in\mathbb{N}$. Then
$\lim_{n\to\infty}\operatorname{D}_{f}^{\mathcal{R}}(\nu_{n}\|\nu)=0.$
Proof.
By (16) and the reverse Fatou lemma,
00 0 $=\int\Bigl{\lvert}f\Bigl{(}\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\Bigr{)}\Bigr{\rvert}\,\mathrm{d}\nu=\int\lim_{n\to\infty}\Bigl{\lvert}f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\Bigr{)}\Bigr{\rvert}\,\mathrm{d}\nu$
$\geq\limsup_{n\to\infty}\int\Bigl{\lvert}f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\Bigr{)}\Bigr{\rvert}\,\mathrm{d}\nu\geq\limsup_{n\to\infty}\Bigl{\lvert}\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu\Bigr{\rvert}\geq 0.$
Hence
$\lim_{n\to\infty}\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu=\lim_{n\to\infty}D^{\mathcal{R}}_{f}(\nu_{n}\|\nu)=0.\qed$
We will need the following “reciprocal” of (16) for the right half of (15).
Proposition 5.3.
Let $\nu$ be a Borel probability measure and $\nu_{n}$ be the corresponding empirical measure, $n\in\mathbb{N}$. Suppose the Radon–Nikodym derivative $\mathrm{d}\nu_{n}/\mathrm{d}\nu$ exists for all $n\in\mathbb{N}$ and are uniformly bounded. Then
$\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu}{\mathrm{d}\nu_{n}}=1.$
Proof.
Let $x_{n}\coloneqq\lvert\operatorname{proj}_{G(\mathcal{R})}(\mathrm{d}\nu/\mathrm{d}\nu_{n})-1\rvert/2$. By the classical Glivenko–Cantelli theorem,
$\lim_{n\to\infty}\int x_{n}\,\mathrm{d}\nu_{n}=\lim_{n\to\infty}\nu_{n}x_{n}=0.$
By assumption, $\mathrm{d}\nu/\mathrm{d}\nu_{n}<c$ for all $n\in\mathbb{N}$ and for some constant $c>0$. If $\lim_{n\to\infty}x_{n}\neq 0$, then there exists $\varepsilon>0$ and a subsequence $(x_{n_{k}})$ such that
$\varepsilon<\int x_{n_{k}}\,\mathrm{d}\nu=\int x_{n_{k}}\frac{\mathrm{d}\nu}{\mathrm{d}\nu_{n_{k}}}\,\mathrm{d}\nu_{n_{k}}\leq c\int x_{n_{k}}\,\mathrm{d}\nu_{n_{k}}\to 0,$
a contradiction, and thus $\lim_{n\to\infty}x_{n}=0$. ∎
Theorem 5.4 (Glivenko–Cantelli theorem for $f$-divergence II).
Let $\mathcal{R}$ be the class of rays, $f\in\mathcal{F}$ as in (14), $\nu$ a Borel probability measure, and $\nu_{n}$ the corresponding empirical measure, $n\in\mathbb{N}$. Then
$\lim_{n\to\infty}\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\nu_{n})=0.$
Proof.
Let $y_{n}\coloneqq\lvert f(\operatorname{proj}_{G(\mathcal{R})}\mathrm{d}\nu/\mathrm{d}\nu_{n})\rvert$. Since $f$ is continuous at $1$,
$\lim_{n\to\infty}y_{n}=\Bigl{\lvert}f\Bigl{(}\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu}{\mathrm{d}\nu_{n}}\Bigr{)}\Bigr{\rvert}=0.$
So
$\Bigl{\lvert}\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu}{\mathrm{d}\nu_{n}}\Bigr{)}\,\mathrm{d}\nu_{n}\Bigr{\rvert}\leq\int\Bigl{\lvert}f\Bigl{(}\operatorname{proj}_{G(\mathcal{R})}\frac{\mathrm{d}\nu}{\mathrm{d}\nu_{n}}\Bigr{)}\Bigr{\rvert}\,\mathrm{d}\nu_{n}=\nu_{n}y_{n}.$
If $\nu_{n}y_{n}\not\to 0$, then there exists $\varepsilon>0$ and a subsequence $(y_{n_{k}})$ such that
$\varepsilon<\lvert\nu_{n_{k}}y_{n_{k}}\rvert\leq\lVert y_{n_{k}}\rVert\to 0,$
a contradiction, and thus $\lim_{n\to\infty}\nu_{n}y_{n}=0$, giving us the required result. ∎
We have established that both $\operatorname{D}_{f}^{\mathcal{R}}(\nu_{n}\|\nu)$ and $\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\nu_{n})$ converge to 00 almost surely. Consequently, either of the following expressions:
$\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)+\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\mu),\quad\text{or}\quad\max\bigl{\{}\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu),\operatorname{D}_{f}^{\mathcal{R}}(\nu\|\mu)\bigr{\}}$
defines a distance that converges to 00 almost surely between the target distribution $\nu$ and the empirical distributions $\nu_{n}$.
6. Discussion
6.1. Choquet integral
In this subsection, we present our initial approach to addressing Question 1 using the Choquet integral [11, 15] and explain why this approach ultimately fails. The Lebesgue integral defines integration with respect to a measure, whereas the Choquet integral extends this concept by defining integration for a more general class of set functions, known as capacities.
Let $\Omega$ be the ground set, and let $\mathcal{S}\subseteq 2^{\Omega}$ be a collection of subsets such that $\varnothing,\Omega\in\mathcal{S}$. A function $\nu:\mathcal{S}\to\mathbb{R}$ is called a capacity on $\mathcal{S}$ if it satisfies the following conditions:
[…]
    Normalization: $\nu(\varnothing)=0$.
[…]
    Monotonicity: For any $A,B\in\mathcal{S}$ with $A\subseteq B$, it holds that $\nu(A)\leq\nu(B)$.
A function $g:\Omega\to\mathbb{R}$ is said to be $\mathcal{S}$-measurable if, for every $t\in\mathbb{R}$, the superlevel set $\{g>t\}$ belongs to $\mathcal{S}$. Denote by $B^{+}(\mathcal{S})$ the set of all bounded, nonnegative, $\mathcal{S}$-measurable functions.
For a function $g\in B^{+}(\mathcal{S})$, the decumulative (distribution) function of $g$ with respect to $\nu$ is given by $G_{\nu,g}(t)=\nu(g>t)$, $t\in\mathbb{R}$. The Choquet integral of $g$ with respect to $\nu$ is then defined as
$\mathsf{C}\!\!\!\!\!\!\int g\,\mathrm{d}\nu=\int_{0}^{\infty}G_{\nu,g}(t)\,\mathrm{d}t,$
where the right-hand side is understood as a Riemann integral.
If the capacity $\nu$ is a measure, then the Choquet integral coincides with the Lebesgue integral [15, p. 62]. A nature approach to defining the $f$-divergence over $\mathcal{R}$ is to set
$\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)\coloneqq\operatorname{D}_{f}(\mu_{\mathcal{R}}\|\nu_{\mathcal{R}}),$
where $\mu_{\mathcal{R}}$ and $\nu_{\mathcal{R}}$ denote the restrictions of $\mu$ and $\nu$ to $\mathcal{R}$.
This idea originates from the observation that it correctly recovers “partial variation” when $\mathcal{R}$ is replaced by a $\sigma$-subalgebra $\mathcal{C}\subseteq\mathcal{B}$. Specifically, if $\mu$ and $\nu$ are probability measures on $\mathcal{B}$, then there restrictions, $\mu_{\mathcal{C}}$ and $\nu_{\mathcal{C}}$, define probability measures on $\mathcal{C}$, leading to:
$\operatorname{D}_{f}^{\mathcal{C}}(\mu\|\nu)\coloneqq\operatorname{D}_{f}(\mu_{\mathcal{C}}\|\nu_{\mathcal{C}})=\sup_{A\in\mathcal{C}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}.$
For the class of rays $\mathcal{R}$, it is straightforward to verify that $\mu_{\mathcal{R}}$ and $\nu_{\mathcal{R}}$ are capacities. Consequently, we initially attempted to define $\operatorname{D}_{f}(\mu_{\mathcal{R}}\|\nu_{\mathcal{R}})$ using Choquet integral, based on the results from [48, 51, 25, 24, 38].
However, a fundamental obstruction arises: there is no guarantee that the Radon–Nikodym derivative $\mathrm{d}\mu_{\mathcal{R}}/\mathrm{d}\nu_{\mathcal{R}}$ exists, nor than the function $f(\mathrm{d}\mu_{\mathcal{R}}/\mathrm{d}\nu_{\mathcal{R}})$, for $f\in\mathcal{F}$, is $\mathcal{R}$-measurable (i.e., nonincreasing) when $\mathcal{R}$ is merely a class of rays rather than a $\sigma$-algebra or even an algebra. As a result, the expression $\operatorname{D}_{f}(\mu_{\mathcal{R}}\|\nu_{\mathcal{R}})$ is not well-defined.
To overcome this issue, we considered an alternative approach by introducing a modified set function:
$\widehat{\mu}_{\mathcal{R}}(A)=\begin{cases}\mu(A),&\text{if }A\in\mathcal{R},\\ 0,&\text{if }A\in\mathcal{B}\setminus\mathcal{R}.\end{cases}$
Similarly, we define $\widehat{\nu}_{\mathcal{R}}$ in the same manner. Then we propose defining:
$\operatorname{D}_{f}^{\mathcal{R}}(\mu\|\nu)\coloneqq\operatorname{D}_{f}(\widehat{\mu}_{\mathcal{R}}\|\widehat{\nu}_{\mathcal{R}}).$
This idea is heuristically motivated by the observation that it correctly recovers the “partial variation distance”. Specifically, for total variation:
$\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{R}}(\mu\|\nu)$ $=\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}(\widehat{\mu}_{\mathcal{R}}\|\widehat{\nu}_{\mathcal{R}})=\sup_{A\in\mathcal{B}}\,\bigl{(}\widehat{\mu}_{\mathcal{R}}(A)-\widehat{\nu}_{\mathcal{R}}(A)\bigr{)}$
$=\max\biggl{\{}\sup_{A\in\mathcal{R}}\,\bigl{(}\widehat{\mu}_{\mathcal{R}}(A)-\widehat{\nu}_{\mathcal{R}}(A)\bigr{)},\sup_{B\in\mathcal{B}\backslash\mathcal{R}}\underbrace{\bigl{(}\widehat{\mu}_{\mathcal{R}}(B)-\widehat{\nu}_{\mathcal{R}}(B)\bigr{)}}_{\equiv 0}\biggr{\}}$
$=\sup_{A\in\mathcal{R}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)},$
However, another fundamental obstacle arises: the set functions $\widehat{\mu}_{\mathcal{R}}$ and $\widehat{\nu}_{\mathcal{R}}$ do not satisfy monotonicity, meaning they are not capacities. Consequently, $\operatorname{D}_{f}(\widehat{\mu}_{\mathcal{R}}\|\widehat{\nu}_{\mathcal{R}})$ cannot be well-defined via the Choquet integral.
6.2. Vapnik–Chervonenkis for $f$-divergence?
In this subsection and the next, we propose two open questions that arise from the results of this paper.
We define a class $\mathcal{C}\subseteq\mathcal{B}$ as a variation class, if for any Borel probability measures $\mu$ and $\nu$, the associated set of functions
$G(\mathcal{C})\coloneqq\bigl{\{}g\in L^{2}(\nu):\{g>r\}\in\mathcal{C},r\in\mathbb{R}\bigr{\}}$
is a closed and convex set in $L^{2}(\nu)$ and the following identity holds:
(17) $\sup_{A\in\mathcal{C}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}=\int\frac{1}{2}\biggl{\lvert}\operatorname{proj}_{G(\mathcal{C})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\biggr{\rvert}\,\mathrm{d}\nu.$
If $G(\mathcal{C})$ is a closed and convex set, we can define the $(\mathcal{C},f)$-divergence as
$\operatorname{D}_{f}^{\mathcal{C}}(\mu\|\nu)=\int f\Bigl{(}\operatorname{proj}_{G(\mathcal{C})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu.$
We impose equation (17) as a necessary condition because we require that the $(\mathcal{C},f)$-divergence satisfies the property
$\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{C}}(\mu\|\nu)=\sup_{A\in\mathcal{C}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}.$
In other words, we want the “$\mathcal{C}$-partial variation distance” to be a special case of $(\mathcal{C},f)$-divergence when $f(t)=\lvert t-1\rvert/2$, similar to how the total variation distance arises as a special case of $f$-divergence. Thus, a natural question follows: for which subcollections $\mathcal{C}\subseteq\mathcal{B}$ is $\mathcal{C}$ a variation class?
We have demonstrated that the class $\mathcal{R}$ is a variation class. If $\mathcal{C}$ satisfies the following conditions:
[…]
    $\{\varnothing,\Omega\}\subseteq\mathcal{C}$, and
[…]
    $\mathcal{C}$ is closed under countable unions and intersections,
then $G(\mathcal{C})$ forms a closed convex cone in $L^{2}(\nu)$.
To see this, note that $G(\mathcal{C})$ is a convex cone because for any $g_{1},g_{2}\in G(\mathcal{C})$ and any nonnegative scalars $c_{1},c_{2}\geq 0$, we have $c_{1}g_{1},c_{2}g_{2}\in G(\mathcal{C})$. Furthermore, the sum of two functions in $G(\mathcal{C})$ remains in $G(\mathcal{C})$ since
$\{g_{1}+g_{2}>c\}=\bigcup_{q\in\mathbb{Q}}\{g_{1}>c-q\}\cap\{g_{2}>q\}\in\mathcal{C}.$
The set $G(\mathcal{C})$ is closed because, if $g_{n}\in G(\mathcal{C})$ is a sequence such that $\lVert g_{n}-g\rVert\to 0$, then $g_{n}\to g$ in $\nu$-measure. Consequently, there exists a subsequence $(g_{\pi(n)})$ such that $g_{\pi(n)}\to g$ almost surely. By the definition of pointwise convergence, we obtain:
$\{g>c\}=\bigcup_{k\geq 1}\bigcap_{m\geq 1}\bigcup_{n\geq m}\Bigl{\{}g_{\pi(n)}>c+\frac{1}{k}\Big{\}}\in\mathcal{C}.$
Since each set on the right-hand side belongs to $G(\mathcal{C})$, and $G(\mathcal{C})$ is closed under countable unions and intersections, it follows that $\{g>c\}\in G(\mathcal{C})$. Hence, $g\in G(\mathcal{C})$, proving that $G(\mathcal{C})$ is closed.
The Borel $\sigma$-algebra $\mathcal{B}$ is a variation class. In particular, when $\mathcal{C}=\mathcal{B}$, we have $G(\mathcal{B})=L^{2}(\nu)$, and thus, for any $g\in L^{2}(\nu)$, the metric projection satisfies $\operatorname{proj}_{G(\mathcal{B})}g=g$. This allows us to recover the standard $f$-divergence as follows:
$\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{B}}(\mu\|\nu)=\int\frac{1}{2}\biggl{\lvert}\operatorname{proj}_{G(\mathcal{B})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\biggr{\rvert}\,\mathrm{d}\nu=\int\frac{1}{2}\biggl{\lvert}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\biggr{\rvert}\,\mathrm{d}\nu=\sup_{A\in\mathcal{B}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}.$
Furthermore, any $\sigma$-subalgebra $\mathcal{C}\subseteq\mathcal{B}$ is also a variation class. In the following, we provide concrete examples to illustrate this property.
Example 6.1.
Let $\mu$ and $\nu$ be Borel probability measures, and let $\mathcal{C}\subseteq\mathcal{B}$ be a $\sigma$-subalgebra. it is clear that $\mathcal{C}$ is a variation class.
Denote by $\mu|_{\mathcal{C}}$ and $\nu|_{\mathcal{C}}$ the restrictions of $\mu$ and $\nu$ to $\mathcal{C}$. Then $\mu|_{\mathcal{C}}$ and $\nu|_{\mathcal{C}}$ remain probability measures [43, p. 30]. Moreover, if $g\geq 0$ is $\mathcal{C}$-measurable, i.e., $\{g>c\}\in\mathcal{C}$ for all $c\in\mathbb{R}$, then for any $A\in\mathcal{C}$, the integral remains unchanged:
$\int_{A}g\,\mathrm{d}\nu|_{\mathcal{C}}=\int_{A}g\,\mathrm{d}\nu.$
Beyond this, if $\mu\ll\nu$, then $\mu|_{\mathcal{C}}\ll\nu|_{\mathcal{C}}$, ensuring that the Radon–Nikodym derivative $\mathrm{d}\mu|_{\mathcal{C}}/\mathrm{d}\nu|_{\mathcal{C}}$ exists. For any $A\in\mathcal{C}$, we have:
$\int_{A}\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}\,\mathrm{d}\nu=\int_{A}\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}\,\mathrm{d}\nu|_{\mathcal{C}}=\mu|_{\mathcal{C}}(A)=\mu(A)=\int_{A}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\,\mathrm{d}\nu.$
This confirms that $\mathrm{d}\mu|_{\mathcal{C}}/\mathrm{d}\nu|_{\mathcal{C}}$ is indeed the conditional expectation, and hence, by [4, p. 90]
(18) $\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}=\mathbb{E}_{\mu}\Bigl{[}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigm{|}\mathcal{C}\Bigr{]}=\operatorname{proj}_{G(\mathcal{C})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}.$
Using this result, we derive:
$\operatorname{D}_{\operatorname{\scriptscriptstyle TV}}^{\mathcal{C}}(\mu\|\nu)$ $=\int\frac{1}{2}\biggl{\lvert}\operatorname{proj}_{G(\mathcal{C})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}-1\biggr{\rvert}\,\mathrm{d}\nu=\int\frac{1}{2}\biggl{\lvert}\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}-1\biggr{\rvert}\,\mathrm{d}\nu$
$=\int\frac{1}{2}\biggl{\lvert}\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}-1\biggr{\rvert}\,\mathrm{d}\nu|_{\mathcal{C}}=\sup_{A\in\mathcal{C}}\,\bigl{(}\mu(A)-\nu(A)\bigr{)}.$
The second open question is: For which variation classes $\mathcal{C}$ does the following hold with probability 1?
$\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{C})}\frac{\mathrm{d}\nu_{n}}{\mathrm{d}\nu}=1,\quad\text{and},\quad\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{C})}\frac{\mathrm{d}\nu}{\mathrm{d}\nu_{n}}=1.$
We refer to such a class $\mathcal{C}$ as a Glivenko–Cantelli class.
The Vapnik–Chervonenkis (VC) theorem [55] implies that if a variation class $\mathcal{C}$ has finite VC dimension, then $\lim_{n\to\infty}\operatorname{proj}_{G(\mathcal{C})}\mathrm{d}\nu_{n}/\mathrm{d}\nu=1$ almost surely. Thus, every variation class with finite VC dimension is a Glivenko–Cantelli class.
However, we will provide an example of a Glivenko–Cantelli class that has an infinite VC dimension, demonstrating that finite VC dimension is sufficient but not necessary for a variation class to satisfy the Glivenko–Cantelli property.
Example 6.2.
Let $\mathcal{E}=(E_{n})_{n\in\mathbb{N}}$ be a countable partition of $\Omega$, and let $\mathcal{C}=\sigma(\mathcal{E})$ be the $\sigma$-algebra generated by $\mathcal{E}$. It follows that any set in $\mathcal{C}$ can be expressed as a countable union of elements in $\mathcal{E}$ [9, p. 4]. Consequently, any $\mathcal{C}$-measurable function $h$ must be a simple function of the form:
$h=\sum_{n=1}^{\infty}\alpha_{n}\mathbbm{1}_{E_{n}}.$
Let $\mu|_{\mathcal{C}}$ and $\nu|_{\mathcal{C}}$ be the restrictions of the probability measures $\mu$ and $\nu$ to $\mathcal{C}$, respectively. The Radon–Nikodym derivative can then be written in closed form as:
$\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}=\sum_{n=1}^{\infty}\alpha_{n}\mathbbm{1}_{E_{n}}.$
For any $m$, we have:
$\mu|_{\mathcal{C}}(E_{m})$ $=\int_{E_{m}}\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}\,\mathrm{d}\nu|_{\mathcal{C}}=\int\mathbbm{1}_{E_{m}}\sum_{n=1}^{\infty}\alpha_{n}\mathbbm{1}_{E_{n}}\,\mathrm{d}\nu|_{\mathcal{C}}$
$=\alpha_{m}\cdot\nu|_{\mathcal{C}}(E_{m}).$
Thus, by Equation (18), the projection of the Radon–Nikodym derivative $\mathrm{d}\mu/\mathrm{d}\nu$ simplifies to:
$\operatorname{proj}_{G(\mathcal{C})}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}=\frac{\mathrm{d}\mu|_{\mathcal{C}}}{\mathrm{d}\nu|_{\mathcal{C}}}=\sum_{n=1}^{\infty}\mathbbm{1}_{E_{n}}\frac{\mu|_{\mathcal{C}}(E_{n})}{\nu|_{\mathcal{C}}(E_{n})}.$
Accordingly, the $(\mathcal{C},f)$-divergence can be expressed as:
$\operatorname{D}_{f}^{\mathcal{C}}(\nu_{n}\|\nu)=\operatorname{D}_{f}(\nu_{n}|_{\mathcal{C}}\|\nu|_{\mathcal{C}})=\sum_{i=1}^{\infty}f\Bigl{(}\frac{\nu_{n}|_{\mathcal{C}}(E_{i})}{\nu|_{\mathcal{C}}(E_{i})}\Bigr{)}\nu|_{\mathcal{C}}(E_{i}).$
Now, assume that the sequence $(\nu_{n}|_{\mathcal{C}}(E_{k})/\nu|_{\mathcal{C}}(E_{k}))_{k}$ is uniformly bounded. Then we obtain:
$\limsup_{n\to\infty}\operatorname{D}_{f}^{\mathcal{C}}(\nu_{n}\|\nu)$ $=\limsup_{n\to\infty}\sum_{i=1}^{\infty}f\Bigl{(}\frac{\nu_{n}|_{\mathcal{C}}(E_{i})}{\nu|_{\mathcal{C}}(E_{i})}\Bigr{)}\nu|_{\mathcal{C}}(E_{i})$
$\leq\sum_{i=1}^{\infty}\limsup_{n\to\infty}f\Bigl{(}\frac{\nu_{n}|_{\mathcal{C}}(E_{i})}{\nu|_{\mathcal{C}}(E_{i})}\Bigr{)}\nu|_{\mathcal{C}}(E_{i})$
$=\sum_{i=1}^{\infty}f\Bigl{(}\lim_{n\to\infty}\frac{\nu_{n}|_{\mathcal{C}}(E_{i})}{\nu|_{\mathcal{C}}(E_{i})}\Bigr{)}\nu|_{\mathcal{C}}(E_{i})$
$=0.$
Thus, we conclude that:
$\lim_{n\to\infty}\operatorname{D}_{f}^{\mathcal{C}}(\nu_{n}\|\nu)=0.$
This result demonstrates that the class $\mathcal{C}=\sigma(\mathcal{E})$, which has infinite VC dimension, still satisfies the Glivenko–Cantelli property.
7. Conclusion
As we noted in Section 1, the total variation distance plays an outsize role in traditional statistics but modern AI often relies on $f$-divergences other than the total variation distance. The work in this article would hopefully serve as a demonstration of how one may extend known results in statistics to cover other $f$-divergences.